### v2

# Elasticity without a reference state: continuum mechanics of active tension nets

Abstract
A constitutive relation between stress and strain relative to a reference state is the basic assumption of elasticity theory. However, in living matter, stress is governed by (motor molecule) activity rather than a constitutive law. What paradigm takes the place of elasticity in this setting? Here, we derive a continuum theory of active mechanics by taking the continuum limit of the Active Tension Network model of 2d epithelia. Instead of a reference state, we start from a prescribed active force configuration, encoded in a Riemannian “tension metric”. Intuitively, one expects cells to adjust their positions to achieve force balance by rearranging local sources of active stress. More precisely, the cell positions define an embedding of the tension metric into 2d physical space, which determines the macroscopic physical stress. For free boundaries, tissue adopts a certain intrinsically defined shape, the force-balanced embedding with minimal internal stress. Boundary forces then deform this embedding. The resulting stress transformation yields an effective stress-strain relation. Key elements of elasticity hence emerge from a “stress-only” starting point, explaining how tissue shape can be adiabatically controlled by active stress during morphogenesis. Plastic behavior arises from topological cell rearrangement, which we represent by a continuous reparameterization of the tension metric, providing a principled continuum theory of emergent elasto-plastic flow. To express this physics, we use the mathematics of isothermal coordinates and quasi-conformal maps. An explicit coarse-graining study, described in a companion paper, substantiates the results of our continuum analysis. The present theory elucidates the unconventional mechanics of living tissues and may apply to 2d active and granular materials more generally.
Introduction
During morphogenesis, cells sculpt the physical form of the developing embryo. In contrast to inert materials, living tissues generate forces internally [1, 2]. Key challenges are to characterize the macroscopic mechanical properties (i.e. the “phase of matter”) of developing tissues and to understand the control of tissue shape (morphogenesis) by the cell-generated forces.
Conventional elasticity, based on a reference shape and a constitutive law relating stress and strain (rate) [3], has served as the starting point for much theoretical work on tissue mechanics [4, 5, 6]. For instance, epithelial cells are often described as having a target area and perimeter [7] with deformation from a target shape defining the stress. Tissue flow on longer time scales is understood in terms of plastic change of the reference state attributed to cell division or rearrangement  [2, 8].
However, biological observations challenge the notions of reference state and constitutive laws. First, the turnover of the force-bearing molecules in cells, rapid compared to morphogenetic timescales, allows the “rest length” of a cell to change [9], ruling out a fixed reference state. Since stress needs to be actively maintained against relaxation due to turnover, one expects that on long timescales, all stress in a tissue is active. Active stress, however, is in principle independent of strain and instead controlled by the activity of molecular motors such as myosin.
We therefore seek to understand tissue mechanics by starting from the specification of cell-associated active stress, instead of a reference shape. Due to the timescale separation between mechanical equilibration (seconds [10]) and morphogenetic remodeling (hours), we expect cells to adjust their shapes and positions to achieve local force balance. Morphogenesis takes the form of an adiabatic transformation between force-balanced states, controlled by changes in the local mechanical activity.
This approach is motivated by the study of confluent, 2D epithelia. Cells in such tissue sheets can regulate their interfacial tensions by molecular motor activity in a cell’s actomyosin cytoskeleton [11, 12], which acts as a “microscopic muscle”. This picture is made precise by the active tension network (ATN) model [11], a minimal model for the mechanics of 2D epithelia. Within this framework, we can ask how cells control tissue shape through local active tensions.
To build intuition, consider a conventional fluid foam. Here, a “cell” (foam bubble) interface has no rest length since its tension is independent of its length, being instead set by the fluid’s surface tension. These surface tensions are balanced against internal pressure in the foam bubbles. Remarkably, even though microscopically fluid, the foam macroscopically behaves like an elastic solid – it can maintain its shape against gravity [13]. In a foam, interfacial tensions are fixed material parameters. In a tissue, they are promoted to active, dynamical variables (under local cellular control), enabling richer behavior.
Here, we derive a continuum theory of emergent, active elasticity. The key to our approach lies in representing the configuration of active stress by a Riemannian metric $\mathbf{g}$, a continuum version of Maxwell–Cremona force tesselations [14]. The cell tiling defines an embedding of the tension metric into physical space. We show that the central elements of elasticity theory, a stress-free reference state and a stress-strain relationship, emerge from this embedding. Additionally, we will demonstrate that cell-scale topological rearrangements that enable large deformations can be captured in the continuum theory via a suitable reparametrization of the tension metric. Taken together, these ideas enable a continuum description of adiabatic tissue flow driven by active stress dynamics. The resulting theory is a form of “emergent” viscoelasticity, in which adiabatic stress dynamics can drive active plastic flow in the absence, or against the resistance, of external forces. In a companion paper [15], we show that the continuum-based, “top-down” approach presented here precisely matches a “bottom-up” coarse-graining analysis of the ATN model.
The remainder of this paper is structured as follows. We begin by reviewing the cell-level ATN model. In Section .1, we explain how we describe cell tiling and tension configuration in the continuum limit. Section .2 studies the static problem, and Section .3 turns to the adiabatic dynamics of active stress. Finally, Sect. .4 considers topological cell rearrangement.
Discrete theory: 2D tension networks in force balance
The microscopic ATN model [11] describes a 2D tiling of cells $i,j,\dots$ whose mechanics is dominated by tensions $\tau_{ij}$ along cell-cell interfaces and intracellular pressure $p_{i}$. In mechanical balance, virtual work has to vanish, $0=dE=-\sum_{ij}\uptau_{ij}d\ell_{ij}+\sum_{i}\mathrm{p}_{i}d\mathrm{A}_{i}$, where $\ell_{ij},\mathrm{A}_{i}$ are interface lengths and cell areas. The pressures $\mathrm{p}_{i}$ are determined by passive bulk elasticity, captured by a local equation of state $\mathrm{p}_{i}=P(\mathrm{A}_{i})$. By contrast, the tensions $\tau_{ij}$ are dynamical variables determined by intrinsic motor molecule activity, and independent of interface length $\ell_{ij}$ – there are no reference lengths and no constitutive laws for tensions. This is the key premise of the ATN model [11, 16].
We aim to understand how the active tensions determine the physical configuration of the cell tiling in force balance, and how this configuration evolves when tensions change adiabatically. To this end, it will be useful to think of the tension configuration as a network dual to the cell tiling. Since the cells generically meet at 3-fold vertices, the tensions form a triangulation with one node per cell $i$ and a length $\uptau_{ij}$ link between neighbors $i$ and $j$; Fig. 1. Each tricellular vertex ijkijk corresponds to one “tension triangle”, the shape of which encodes the local tension configuration (for example, tension anisotropy).
Results
.1 Continuum description of ATNs as Riemannian surfaces
.1.1 Tension metric and macroscopic stress tensor
A triangulation, specified by its adjacency graph and edge lengths, defines a (piecewise-linear) metric surface  [^1] The definition of the manifold structure is explained in detail in our companion paper [15]. (see Fig. 1, top). We can therefore coarse-grain the tension triangulation to a continuous tension manifold with tension metric $g_{\alpha\beta}(\bm{\xi})$ as a function of Lagrangian coordinates $\bm{\xi}$, which play the role of the cell labels $i,j$ For an infinitesimal line element $d\bm{\xi}$, the tension metric yields the tension $d\tau$ on cell interfaces oriented transverse to it as $d\tau^{2}=g_{\alpha\beta}(\bm{\xi})d\xi_{\alpha}d\xi_{\beta}$. In the “microscopic limit” where $d\bm{\xi}$ connects two adjacent triangulation nodes, $d\tau$ is the tension along the interface between the corresponding cells.
Figure 1: A triangulation (left) is a discrete Riemannian surface. The tension metric defines junctional tensions. An embedding of the metric into physical space is the continuum equivalent of the centroids in the cell tessellation.
Note that different triangulations (i.e., local tension configurations) can realize the same tension manifold. For instance, all triangulations of the plane realize the Euclidean metric. The tension triangulation thus encodes “microscopic” information in addition to the tension metric $\mathbf{g}$; we return to this point systematically in Sect. .4. One important additional degree of freedom is the local cell density $n({\bm{\xi}})d^{2}\xi$, the number of cells (triangulation nodes) per unit area.
A set of physical cell positions $\mathbf{r}({\bm{\xi}})$ now defines a map from the tension manifold into two-dimensional physical space, to which we refer as “embedding” (Fig. 1, bottom). For simplicity, we assume that physical space is the plane, but our theory readily generalises to tissues on arbitrary curved surfaces. Under the map $\mathbf{r}({\bm{\xi}})$, the metric $\mathbf{g}$ transforms as
$g_{\alpha\beta}(\mathbf{r})=\frac{\partial\xi_{\gamma}}{\partial r_{\alpha}}g_{\gamma\delta}(\bm{\xi})\frac{\partial\xi_{\delta}}{\partial r_{\beta}}$ (1)
In the following, we non-dimensionalize $\mathbf{g}\mapsto p_{0}^{2}\mathbf{g}$ so the metric has units of $\mathrm{length}]^{-2}$. The dimensional factor $p_{0}$ converts units of line tension to units of length, and will be identified as the reference pressure below.
We now link the geometry of the tension metric to the mechanics of the cell. The macroscopic stress tensor of an ATN is the sum of two contributions
$\sigma^{\mathrm{tot}}_{\alpha\beta}=\sigma_{\alpha\beta}-p\delta_{\alpha\beta}$ (2)
where $\sigma_{\alpha\beta}$ is the tensional stress due to tensions $\uptau_{ij}$, and $p$ the intracellular pressure. The original ATN model [11] assumed uniform intracellular pressure. Here, we generalize: intracellular pressure is determined by an equation of state $p=P(n)$ as a function of cell density in the continuum. The reference cell density $n_{0}$ is defined by $P(n_{0})=p_{0}$. (In contrast to the active tensile stress $\sigma$, we treat pressure as a passive field, hence the equation of state.) Under $\mathbf{r}({\bm{\xi}})$, the density transforms as $n(\mathbf{r})=n({\bm{\xi}})\,{\det[\partial_{\bm{\xi}}\mathbf{r}]}$, and therefore:
${}p(\mathbf{r})=P\bigl(n({\bm{\xi}})\det[\partial_{\bm{\xi}}\mathbf{r}]\bigr)$ (3)
In the following, we refer to $p$ simply as pressure, noting that it differs from the total pressure $\mathrm{Tr}[\bm{\sigma}^{\mathrm{tot}}]/2=p-\mathrm{Tr}[\bm{\sigma}]/2$.
The tensile stress $\bm{\sigma}$ is determined by the tension metric. Consider two adjacent cells at $\mathbf{r}$ and $\mathbf{r}+d\mathbf{r}$ (we therefore implicitly set $dr$ to be of the scale of a single cell, which becomes infinitesimally small in the continuum limit). The magnitude of the force $d\bm{f}$ acting through the virtual cut $d\mathbf{r}$ (Fig. 2) equals the tension $|d\bm{f}|=d\tau$ along the interface between the two cells. It is determined by the tension metric, $d\tau^{2}=p_{0}^{2}\,g_{\alpha\beta}(\mathbf{r})dr_{\alpha}dr_{\beta}$. On the other hand, the stress tensor applied to the cut normal yields $d\bm{f}=\sigma\cdot\mathbf{n}\cdot dr=\mathbf{\sigma}\cdot\epsilon\cdot d\mathbf{r}$, where $\epsilon$ is the antisymmetric tensor, $\mathbf{n}$ is the unit normal, and $\cdot$ is the standard inner product in Euclidean physical space. Equating the tension- and cell-space expressions:
$p_{0}^{2}d\mathbf{r}^{T}\cdot\mathbf{g}(\mathbf{r})\cdot d\mathbf{r}=d\tau^{2}=d\mathbf{r}^{T}\cdot(\epsilon^{T}\cdot\bm{\sigma}^{2}(\mathbf{r})\cdot\epsilon)\cdot d\mathbf{r}$ (4)
Since this equation must hold for arbitrary $d\mathbf{r}$, the stress tensor must be the matrix square root of the metric, rotated by $\pi/2$, and scaled by $p_{0}$. This relation between microscopic stress, encoded in the tension metric, and macroscopic stress is a central element of our continuum theory. Eq. (4) and the stress tensor are defined in physical space, not on the tension manifold, and depend on the cell positions $\mathbf{r}({\bm{\xi}})$, which are physical observables, not mere coordinates.
In mechanical equilibrium, forces must balance:
$\partial_{\alpha}\sigma_{\alpha\beta}^{\mathrm{tot}}=\partial_{\alpha}\sigma_{\alpha\beta}-\partial_{\beta}p=0$ (5)
Equations (4) and (5), supplemented with mechanical boundary conditions (BCs), set the physical configuration (cell positions) $\mathbf{r}({\bm{\xi}})$ for a given tension metric $\mathbf{g}({\bm{\xi}})$ and cell density $n({\bm{\xi}})$. In the following, we introduce a mathematical framework to solve Eqs. (4)–(5). We first study the statics problem to see how tissue shape and response to external forces emerge from (fixed) $\mathbf{g}({\bm{\xi}})$. Then, we turn to quasi-static shape dynamics due to changes of $\mathbf{g}({\bm{\xi}})$, and finally generalize our framework to describe plastic flow (cell rearrangement).
Figure 2: Left: In the continuum limit, the displacement vector between two cell centroids becomes a differential $d\mathbf{r}$. The tension metric determines the tension $d\tau$ on the interface between these cells. In isothermal coordinates, $dr=p_{0}^{-1}d\tau$. Right: The stress tensor $\mathbf{\sigma}$ is defined by the traction forces that act through infinitesimal virtual cuts. A cut along $d\mathbf{r}$ must yield the interfacial tension, relating tension metric and stress tensor in Eq. (4).
.1.2 Complex coordinates and (quasi)conformal maps
To solve Eq. (5), we use complex coordinates $z=z_{1}+iz_{2}$. A symmetric matrix $\mathbf{q}$ (e.g. the metric $\mathbf{g}$) defines a quadratic form $\mathbf{v}\mapsto q_{\alpha\beta}v^{\alpha}v^{\beta}$ which becomes
$4q_{\alpha\beta}v^{\alpha}v^{\beta}$ $=(q_{z\bar{z}}+q_{\bar{z}z})v\bar{v}+q_{zz}vv+q_{\bar{z}\bar{z}}\bar{v}\bar{v},$ (6a)
$\mathrm{where}\;\;q_{z\bar{z}}$ $=q_{\bar{z}z}=q_{11}+q_{22}=\mathrm{Tr}\,\mathbf{q},$
$q_{zz}$ $=\overline{q_{\bar{z}\bar{z}}}=q_{11}-q_{22}+2iq_{12}$ (6b)
Since $q_{zz}$ is the deviatoric component of $\mathbf{q}$, the Beltrami coefficient $\mu_{q}$ measures the relative anisotropy
$\mu_{q}=q_{zz}/\mathrm{Tr}[\sqrt{\mathbf{q}}]^{2}$ (7)
Note that as a coordinate, the complex number $z=z_{1}+iz_{2}$ contains the same information as the Cartesian vector $(z_{\alpha})=\mathbf{z}=(z_{1},z_{2})$. We use the two interchangeably. However, it is convenient to write functions $f$ as $f(z,\bar{z})$ with formally independent variables $z$, $\bar{z}$. With this convention, vector calculus can be expressed in terms of the Wirtinger derivatives $\partial_{z}=(\partial_{1}-i\partial_{2})/2,\;\bar{\partial}_{z}=(\partial_{1}+i\partial_{2})/2$. For example, for real-valued potentials $\theta,\psi$, $2\bar{\partial}_{z}\theta$ is the gradient, while $2i\bar{\partial}_{z}\psi$ is a solenoidal vector field. The Euclidean Laplace operator is $4\partial_{z}\bar{\partial}_{z}=\partial_{1}^{2}+\partial_{2}^{2}=:\Delta$. The Jacobian matrix $\mathbf{D}\kern-0.80002ptf$} of a map $f(z,\bar{z})$ reads
${\mathbf{D}\kern-0.80002ptf}=\begin{bmatrix}\partial_{z}f&\bar{\partial}_{z}f\\[3.99994pt] \partial_{z}\bar{f}&\overline{\partial_{z}f}\end{bmatrix}$ (8)
Note that the matrix entries in Eq. (8) are not in Cartesian $z_{1},z_{2}$ but in complexified $z,\bar{z}$ coordinates.
Maps satisfying the Cauchy–Riemann equation $\bar{\partial}_{z}f=0$, i.e. $z\mapsto z^{\prime}=f(z)$, are conformal (angle-preserving). Their Jacobian is a diagonal matrix ${\mathbf{D}\kern-0.80002ptf}=\mathrm{diag}[\partial_{z}f,\overline{\partial_{z}f}]$, combining rotation and scaling by the conformal factor $\lambda_{z|z^{\prime}}=|\partial_{z}f|$. Furthermore, in this case, $\Delta\log\lambda_{z|z^{\prime}}=4\partial_{z}\bar{\partial}_{z}\operatorname{Re}[\log(\partial_{z}f)]=0$. More generally, the shear created by a map is measured by its Beltrami coefficient $\mu_{z|z^{\prime}}=\bar{\partial}_{z}f/\partial_{z}f$ (the subscript of $\mu_{z|z^{\prime}}$ denotes the map’s source and target.) Note that the Beltrami coefficients of the map $f$ and the quadratic form $({\mathbf{D}\kern-0.80002ptf})^{\dagger}\cdot{\mathbf{D}\kern-0.80002ptf}$ agree; $()^{\dagger}$ being the Hermitian adjoint. Geometrically, $f$ maps infinitesimal circles to ellipses with eccentricity $|\mu_{z|z^{\prime}}|/(1-|\mu_{z|z^{\prime}}|^{2})^{1/2}$. If $|\mu_{z|z^{\prime}}|<1$ everywhere, $f$ is called quasi-conformal (QC). In particular, QC maps preserve orientation. Notation-wise, their argument distinguishes QC maps $f(z,\bar{z})$ (also written $f(\mathbf{z})$), from conformal maps $f(z)$. In the following, we use the formalism of QC maps to develop our continuum theory (in Ref. [18], a similar QC approach was used to describe the growth of 2d tissues).
.1.3 Isothermal coordinates
We now use the mathematics of QC maps to define a convenient parametrization of the tension triangulation. In 2D, a metric has 3 independent components, while an embedding map has two degrees of freedom. Thus, only a scalar degree of freedom remains. It is therefore always possible to find a set of isothermal coordinates [19] in which the metric is isotropic, characterized by a single scale factor. We begin by identifying isothermal coordinates $z(\xi,\bar{\xi})$ for the tension metric. Using Eq. (6), isotropy requires
$4g_{\alpha\beta}d\xi^{a}d\xi^{b}$ $=\mathrm{Tr}[\sqrt{\mathbf{g}}]^{2}|d\xi+\mu_{g}d\bar{\xi}|^{2}$ (9)
$\stackrel{{\scriptstyle!}}{{=}}2g_{z\bar{z}}dzd\bar{z}=g_{z\bar{z}}|\partial_{\xi}zd\xi+\bar{\partial}_{\xi}zd\bar{\xi}|^{2}$
where the metric’s Beltrami coefficient $\mu_{g}=g_{zz}/\mathrm{Tr}[\sqrt{\mathbf{g}}]^{2}$ measures the anisotropy of $\mathbf{g}$. Therefore, $z(\xi,\bar{\xi})$ must fulfill a Beltrami equation:
$\bar{\partial}_{\xi}z(\xi,\bar{\xi})=\mu_{g}\partial_{\xi}z(\xi,\bar{\xi})$ (10)
(Equivalently, $\mu_{\xi|z}=\mu_{g}$.) The QC map from Lagrangian $\xi$- to isothermal $z$-coordinates is exactly such that it removes the shear part of the metric. (By abuse of notation, we will denote both the coordinates and the map $\xi\mapsto z(\xi,\bar{\xi})$ by $z$). The metric, therefore, becomes isotropic and is fully described by its conformal factor $\lambda_{g}$
$g_{zz}=0,\;g_{z\bar{z}}=2\lambda_{g}^{2}(z,\bar{z})\;\Leftrightarrow\;g_{\alpha\beta}=\lambda_{g}^{2}\delta_{\alpha\beta}$ (11)
Isothermal coordinates are conformal and do not distort angles (see Fig. 3). Appendix G.2 discusses an example (isothermal coordinates for a hemisphere).
In isothermal coordinates, the Liouville equation determines the Gaussian curvature:
$K_{g}=-\lambda_{g}^{-2}\Delta\log\lambda_{g}$ (12)
where $\Delta=4\partial_{z}{\bar{\partial}}_{z}$ is the Euclidean Laplacian. More generally, the St. Venant formula [20] gives the curvature of $g_{\alpha\beta}=\lambda_{g}^{2}(\delta_{\alpha\beta}+dg_{\alpha\beta})$ to linear order:
${}K_{g}=-\lambda_{g}^{-2}\big($ $\Delta\log\lambda_{g}+2\operatorname{Re}[\partial_{z}\bar{\partial}_{z}dg_{\bar{z}z}+\partial_{z}^{2}dg_{zz}]\big)$ (13)
Note that the composition of a solution to Eq. (10) with any conformal map yields another solution (indeed, for conformal maps, $\Delta\log\lambda_{z|z^{\prime}}=0$, so Eq. (12) is invariant). The BCs fix the conformal (“homogeneous”) solution. Throughout, we use “natural BCs”
$\lambda_{g}(z,\bar{z})|_{\partial\Omega}=\big(\lambda_{g}(\xi,\bar{\xi})|\partial_{\xi}z|\big)\big|_{\partial\Omega}=1$ (14)
where $\partial\Omega$ is the system boundary. If $\mathbf{g}$ is flat, this BC ensures $g_{z\bar{z}}=2$, $\lambda_{g}=1$ everywhere (Fig. 3).
Figure 3: Isothermal coordinates “flatten” the tension manifold without shear. We isothermally map to the “natural domain” defined by the boundary condition $\lambda_{g}|_{\partial\Omega}=1$. The quasi-conformal map $w(z,\bar{z})$ maps the reference configuration to the physical domain. It determines the macroscopic stress tensor $\bm{\sigma}$ and is subject to physical boundary conditions (gray shading illustrates a clamped boundary, red arrows illustrate a traction force along the boundary).
.2 Statics: Emergent elasticity
We introduced isothermal $z(\xi,\bar{\xi})$ as a convenient parametrization of the tension manifold. We now interpret $z(\xi,\bar{\xi})$ as a specific physical configuration, i.e. a map that places the Lagrangian material element $\bm{\xi}$ at physical position $\mathbf{r}(\bm{\xi})=\mathbf{z}(\bm{\xi})$ in the plane.
.2.1 Emergent reference state
In the $\mathbf{g}$-isothermal configuration, the tensile stress becomes isotropic. Indeed, by Eq. (4):
$\sigma_{zz}=0,\;\sigma_{z\bar{z}}=2p_{0}\lambda_{g}(z,\bar{z})$ (15)
Therefore, with a pressure field
$p=p_{0}\lambda_{g}(z,\bar{z}),$ (16)
the macroscopic stress can be made to vanish identically: $\sigma_{\alpha\beta}^{\text{tot}}=\sigma_{\alpha\beta}-p\delta_{\alpha\beta}=0$. The $\mathbf{g}$-isothermal configuration, thus, is a stress-free reference configuration which emerges from a stress-only starting point. A liquid bubble raft provides a simple physical example: Surface tension balances against pressure so that the raft as a whole is stress-free (cutting it does not lead to macroscopic recoil). Equations (16) and (12) together imply a Laplace equation for the pressure in the stress-free configuration: $\Delta_{z}\log(p/p_{0})=-K_{g}$.
In a physical configuration, the pressure field $p$ is not an independent variable, but must obey the equation of state $p=P(n)$, Eq. (3). The stress-free reference state is physically attainable only if $p_{0}\lambda_{g}(\mathbf{z})=P(n(\mathbf{z}))$ where $n(\mathbf{z})$. For a flat tension metric, this relation reduces to
$n(\mathbf{z})=n({\bm{\xi}})\det[\partial_{\bm{\xi}}\mathbf{z}]\stackrel{{\scriptstyle?}}{{=}}n_{0}$ (17)
Eq. (17) does not hold in general. If it is not fulfilled, the $\mathbf{z}$-configuration is not mechanically balanced, and the tissue adopts a different configuration, in general, with residual stresses  [^2] Residual stresses are those present even for traction-free BCs. Note that the “incompatibility” of the tension manifold with the cell density that gives rise to residual stresses is distinct from the geometric incompatibility of a curved “reference metric” with a flat physical configuration found in the metric elasticity framework [52]. In the latter, the reference metric encodes local rest lengths in the material.. Nonetheless, the isothermal $\mathbf{z}$-configuration will be a useful reference to understand the resulting general force-balanced states. In particular, it will be the starting point for linearizing the theory.
.2.2 General force balanced states and conformal-isogonal decomposition
To relate the stress-free reference state $\mathbf{z}({\bm{\xi}})$ to the realized physical state, we need to calculate how the macroscopic tensile stress $\bm{\sigma}$ changes when the system is deformed from the isothermal reference into a new configuration, $z\mapsto w(z,\bar{z})$. Physically, a deformation moves and reorients the local active force dipoles (cell interfaces), thereby modifying the macroscopic stress tensor. Mathematically, $\mathbf{g}$ transforms according to Eq. (1). Using $g_{ab}(\mathbf{z})=\lambda_{g}(\mathbf{z})\delta_{ab}$, Eq. (4) becomes:
$[\bm{\sigma}^{2}(\mathbf{w})]_{\alpha\beta}$ $=p_{0}^{2}\epsilon_{\alpha\gamma}g_{\gamma\delta}(\mathbf{w})\epsilon_{\beta\delta}=p_{0}^{2}\lambda_{g}^{2}(\mathbf{z})\epsilon_{\alpha\gamma}\frac{\partial z_{\kappa}}{\partial w_{\gamma}}\frac{\partial z_{\kappa}}{\partial w_{\delta}}\epsilon_{\beta\delta}$
The stress is thus determined by the square root of the Jacobian $\partial z^{\alpha}/\partial{w_{\beta}}$. In complexified notation:
$\begin{bmatrix}\sigma_{{\bar{w}}w}&\sigma_{{\bar{w}}{\bar{w}}}\\ \sigma_{ww}&\sigma_{w{\bar{w}}}\end{bmatrix}=2p_{0}\lambda_{g}\epsilon^{T}[\mathbf{D}w^{\dagger}\mathbf{D}w]^{-\frac{1}{2}}\epsilon$ (18)
and the inverse Jacobian $(\mathbf{D}w)^{-1}$ reads
$\hskip-5.0pt\mathbf{D}w^{-1}=\begin{bmatrix}\partial_{w}z&\bar{\partial}_{w}z\\ \partial_{w}\bar{z}&\overline{\partial_{w}z}\end{bmatrix}=\begin{bmatrix}\partial_{w}z&0\\ 0&\overline{\partial_{w}z}\end{bmatrix}\cdot\begin{bmatrix}1&\mu_{w|z}\\ \overline{\mu_{w|z}}&1\end{bmatrix}$ (19)
where $\mu_{w|z}=\bar{\partial}_{w}z/\partial_{w}z$ is the Beltrami coefficient $w(z,\bar{z})$. The $\mu$-matrix factor is hermitian, which makes taking the square root in Eq (18) trivial. The resulting stress tensor components read
$\sigma_{w{\bar{w}}}=2p_{0}\lambda_{g}|\partial_{w}z|,\quad\sigma_{ww}=-2p_{0}\lambda_{g}|\partial_{w}z|\mu_{w|z}$ (20)
Eq. (20) relates spatial deformation to the changes in physical stress, and can be interpreted as an emergent stress-strain relationship. The anisotropic strain relative to the isothermal reference state sets the deviatoric stress $\sigma_{ww}$. Crucially, the stress anisotropy is not identical to the “microscopic” tension anisotropy (differences in interface tensions $\tau_{ij}$, or, geometrically, anisotropy of tension triangles). We will return to this important distinction in Sect. .4 where we define a tension anisotropy field in the continuum.
In Cartesian notation, the matrix square-root Eq. (18) can be evaluated via the singular value decomposition of $\mathbf{D}w=U\cdot\Lambda\cdot V^{T}$:
$\bm{\sigma}(\mathbf{w})=p_{0}\lambda_{g}\frac{U\cdot\Lambda\cdot U^{T}}{\det\Lambda}$ (21)
$\bm{\sigma}$ is therefore symmetric and positive semi-definite, as required for a contractile/tensile stress.
However, not all deformations can lead to a mechanically balanced state. In complexified notation, force balance reads
$\bar{\partial}_{w}\sigma_{{\bar{w}}w}+\partial_{w}\sigma_{ww}$ $=2\bar{\partial}_{w}p$ (22)
As $\partial_{\alpha}\sigma_{\alpha\beta}=\partial_{\beta}p$, the net tensile force $\partial_{\alpha}\sigma_{\alpha\beta}$ must be curl-free:
$\partial_{w}\bar{\partial}_{w}\sigma_{{\bar{w}}w}+\partial_{w}^{2}\sigma_{ww}$ $=2\partial_{w}\bar{\partial}_{w}p$ (23)
The RHS is real, which entails a solvability condition:
$\operatorname{Im}\!\left[\partial_{w}^{2}\sigma_{ww}\right]=\operatorname{Im}\!\left[\partial_{w}^{2}\left(\lambda_{g}\frac{|\partial_{w}z|}{\partial_{w}z}\bar{\partial}_{w}z\right)\right]=0.$ (24)
For a flat tension metric ($K_{g}=0$), one has $\lambda_{g}=1$, so this condition can be further simplified. In Cartesian coordinates (further assuming small displacement $|w-z|\ll 1$):
$\Delta(\epsilon_{\alpha\beta}\partial_{\beta}w_{\alpha}(\mathbf{z}))=0$ (25)
That is, the curl of the deformation $w$ must be harmonic. Equation (24) constrains force-balanced physical configurations $w(z,\bar{z})$. Two classes of solutions are readily apparent.
Conformal deformations.
First, conformal maps $z=f(w)$, for which $\partial_{w}\bar{z}=\overline{\partial_{\bar{w}}z}=0$, so Eq. (24) holds trivially. Under conformal maps, $\sigma$ remains isotropic, while the pressure required for force balance transforms like $p\mapsto p/\lambda_{z|w}$. This allows broadening the conditions under which a stress-free configuration exists. If the reference state $\mathbf{z}({\bm{\xi}})$ fails to respect $n(\mathbf{z})=n_{0}$ (Eq. (17)), we can find a conformal map $w(z)$ so that $n(\mathbf{w})=\lambda_{z|w}^{-2}n(\mathbf{z})=n_{0}$, as long as $\Delta\log n(\mathbf{z})=0$ (which restriction is due to $\Delta\log\lambda_{z|w}=0$ for conformal maps).
As an aside, we note that the relation $p=p_{0}/\lambda_{z|w}$ between conformal deformations, which curve line elements, and pressure gradients can be understood as a continuum manifestation of the Young–Laplace law [13]. Consider a straight line with normal $\mathbf{n}$ in the reference configuration ($z$-coordinates). Through a conformal deformation it acquires a curvature $1/R=\mathbf{n}\cdot(2\bar{\partial}_{w}|\partial_{w}z|)$ (Appendix A), while the pressure drop $dp$ across the line element changes to $p_{0}\mathbf{n}\cdot(2\bar{\partial}_{w}|\partial_{w}z|)dr$. Recall that in $\mathbf{g}$-isothermal coordinates, $p_{0}dr=d\tau$ is the effective line tension, so that $d\tau/R=dp$ – a differential form of the Young–Laplace law.
Isogonal deformations.
Second, for $\lambda_{g}=1$, a map $z(w,\bar{w})=\bar{\partial}_{w}\theta$ defined by the gradient of a real potential $\theta$ is a solution since
$\operatorname{Im}\!\left[\partial_{w}^{2}\left(\tfrac{|\partial_{w}z|}{\partial_{w}z}\bar{\partial}_{w}z\right)\right]=\operatorname{Im}\!\left[\partial_{w}^{2}\bar{\partial}_{w}^{2}\theta\right]=0$ (26)
In contrast to conformal maps, potential gradients generically create anisotropic stress. Indeed, by Eq. (20), $\theta$ is the Airy function of the resulting stress:
$\sigma_{w\bar{w}}=-8p_{0}\,\partial_{w}\bar{\partial}_{w}\theta,\;\sigma_{ww}=8p_{0}\,\bar{\partial}_{w}^{2}\theta$ (27)
An analogous, if lengthy, calculation shows that for $K_{g}\neq 0$, $z=\lambda_{g}^{-2}\bar{\partial}_{w}\theta$ is a perturbative solution to Eq. (24) (it corresponds to the Riemannian gradient of $\theta$ with respect to $\mathbf{g}$). The manifold of force-balanced states parametrized by $\theta$ corresponds to the “isogonal” modes of the discrete ATN model [22, 11], and we will therefore refer to $\theta$ as the isogonal potential (this connection is detailed in the companion paper [15]). Conformal and isogonal displacements exhaust all solutions to Eq. (24) (Appendix B). Any physical configuration $w$ can hence be obtained from the reference $z$ by a composition of a potential map and a conformal map. The two modes suffice to accommodate arbitrary deformations imposed on the system boundary (Appendix C).
In summary, starting from Lagrangian labels $\xi$, we mapped first to the $\mathbf{g}$-isothermal reference configuration $z$, and then to the physical configuration $w$:
$\xi\xmapsto{\mathbf{g}-\mathrm{isothermal}}z(\xi,\bar{\xi})\xmapsto{\text{force balance}}w(z,\bar{z}).$ (28)
Appendix G.1 provides an elementary example of this mapping. While force balance in the bulk imposes a conformal–isogonal structure on the deformation $w$, BCs and a specific constitutive relation for pressure are required to fully determine the physical configuration.
.2.3 Fully compressible cells
By Eq. (24), for all isogonal–conformal configurations, there is a pressure field that brings the system into mechanical equilibrium. However, physically, the pressure field must also obey the equation of state $P(n)$, which we impose now.
Let us first discuss a specific case, the continuum equivalent to the “classic ATN” setting investigated in  [11]. For a flat tension metric $\mathbf{g}$, Eq. 16 implies a constant pressure $p=p_{0}$. In the bulk, tensions are in force balance “on their own”: only at the boundary, a pressure drop $p_{0}$ is required for normal force balance. This state can be realized by freely compressible cells with an “equation of state” $p=p_{0}$ (i.e., cells can freely exchange volume with one another, and $p_{0}$ acts as a Lagrange multiplier for the total tissue area).
Purely isogonal deformations of the reference state also do not require pressure gradients, since the tensile stress is balanced on its own:
$\hskip-3.99994pt\bar{\partial}_{w}\sigma_{w\bar{w}}+\partial_{w}\sigma_{ww}=8p_{0}\,[\bar{\partial}_{w}(\partial_{w}\bar{\partial}_{w}\theta)-\partial_{w}(\bar{\partial}_{w}^{2}\theta)]=0$ (29)
Since the equation of state is trivial, $P(n)=p_{0}$, pressure also remains invariant. Isogonal modes in the bulk are therefore true soft modes (costing zero energy) in this case.
Under conformal deformations $w(z)$, however, the isotropic stress $\sigma_{w\bar{w}}=2p_{0}|\partial_{w}z|$ becomes non-uniform, implying pressure gradients. To sustain such pressure gradients, a finite compressibility (a non-constant $P(n)$) is required. Now, force balance $\partial_{\beta}p=\partial_{\alpha}\sigma_{\alpha\beta}$ selects a single physical solution among the isogonal–conformal family of solutions by fixing the isogonal potential $\theta$ in the bulk. Mechanical BCs further fix the boundary DOFs of the conformal and isogonal maps.
.2.4 Linearized theory: effective Cauchy elasticity
To understand the physics resulting from the interplay of conformal–isogonal structure and the constitutive law for pressure, we now expand stress and pressure for small deformations $w=z+u(z,\bar{z}),\;|\partial_{z}u|\ll 1$ around the isothermal, stress-free reference state. We further assume small cell-density gradients in the reference state, $n(z,\bar{z})=n_{0}(1+\delta n(z,\bar{z}))$, and small curvature $K_{g}$ (hence, $\log\lambda_{g}$ is also small). In the discrete setting, cell areas are close to the Voronoi areas of the tension triangulation.
Near the reference cell density $n_{0}$, one can linearize the equation of state $P(n)\approx p_{0}(1+B(n/n_{0}-1))$, where $B$ is the non-dimensionalized cellular bulk modulus. Evaluating at $n(w,\bar{w})$, one obtains
$p(w,\bar{w})=p_{0}\big\{1-B(\operatorname{Re}[\partial_{z}u]+\delta n)\big\}$ (30)
Also expanding the tensile stress Eq. (20) to linear order in $u$, the total stress becomes
$\sigma^{\mathrm{tot}}_{w\bar{w}}$ $=2p_{0}\big\{(2B-1)\operatorname{Re}[\partial_{z}u]+\log\lambda_{g}-2B\delta n\big\}$
$\sigma^{\mathrm{tot}}_{ww}$ $=2p_{0}\partial_{z}\bar{u}$ (31)
In Cartesian notation, one finds
$\sigma_{\gamma\gamma}^{\mathrm{tot}}$ $=2p_{0}\left(B-\tfrac{1}{2}\right)\partial_{\gamma}u_{\gamma}+2p_{0}(\log\lambda_{g}+B\delta n)$
$\sigma_{\alpha\beta}^{\mathrm{tot}}-\tfrac{1}{2}\delta_{\alpha\beta}\sigma_{\gamma\gamma}^{\mathrm{tot}}$ $=\frac{p_{0}}{2}(\partial_{\alpha}u_{\beta}+\partial_{\beta}u_{\alpha}-\delta_{\alpha\beta}\partial_{\gamma}u_{\gamma})$ (32)
Eq. (.2.4) is mathematically equivalent to linear isotropic elasticity [3] (even though the underlying physical picture is quite different). One identifies the effective shear and bulk moduli as $\frac{1}{2}p_{0}$ and $(B-1/2)p_{0}$. At higher orders in the gradient expansion, the elastic moduli scale like $\sim p$ and thus become non-uniform (high-pressure regions are stiffer). The finite shear modulus indicates that active tension networks are effective solids, rather than fluids. Furthermore, for a sufficiently large cellular bulk modulus $B$, both effective moduli are positive. The constitutive relation $P(n)$ thus ensures that the mechanical equilibrium is stable. If pressures are instead upgraded to independent, active degrees of freedom (like the tensions $\uptau_{ij}$), mechanical stability is no longer guaranteed. For example, fixing both pressure and surface tension in a soap bubble renders it unstable.
By Eq. (.2.4), force balance $\partial_{\alpha}\sigma^{\mathrm{tot}}_{\alpha\beta}=0$ becomes:
$\Delta u_{\beta}+\left(2B-1\right)\partial_{\beta}$ $(\partial_{\alpha}u_{\alpha})=-2\partial_{\beta}\left(B\delta n+\log\lambda_{g}\right)$ (33)
As expected from the solvability analysis above, the general solution to Eq. (33) can be written as $u_{\alpha}=\partial_{\alpha}\theta+\nu_{\alpha}$ where $\nu_{\alpha}$ is a conformal vector field [23, 20]. Eq. (33) reduces to:
$\Delta\theta=-(1-(2B)^{-1})\;\partial_{\alpha}\nu_{\alpha}-\delta n-B^{-1}\log\lambda_{g}+C$ (34)
where $C$ is an integration constant which can be absorbed into a uniform scaling. Using Eq. (34) and the Cauchy–Riemann equations for $\nu_{\alpha}$, the stress can be brought into the form:
$\sigma_{\alpha\beta}^{\mathrm{tot}}=p_{0}\left(\partial_{\alpha}\partial_{\beta}\theta-\Delta^{2}\theta\delta_{\alpha\beta}\right)=-p_{0}\,\epsilon_{\alpha\gamma}\epsilon_{\beta\delta}\partial_{\gamma}\partial_{\delta}\theta$ (35)
Note that Eq. (35) is valid only in a force-balanced state, i.e. when Eq. (34) holds. For an arbitrary, unbalanced deformation $u_{\alpha}$, one must use Eq. (.2.4).
Hence, the isogonal potential $\theta$ is identical to the Airy stress function  [^3] More precisely, it is the potential of the inverse mapping $z(w,\bar{w})=w+2\bar{\partial}_{w}\theta$ which defines the Airy function. To linear order, this only changes a sign.. Due to Eqs. (12) and (34), $\theta$ obeys a biharmonic equation analogous to classical elasticity:
$\Delta^{2}\theta=\Delta\delta n-B^{-1}K_{g}$ (36)
Gaussian curvature of the tension metric and $\delta n$ appear as source terms, implying residual stresses even for free BCs. These stresses result from an incompatibility between a constant cell density/pressure and an isotropic tensile stress (the macroscopically stress-free state does not have uniform pressure). The biharmonic equation Eq. (36), together with BCs on stress $\sigma_{\alpha\beta}$ or displacement $u_{\alpha}$, fully determines the tissue shape. It can be solved numerically or via complex variable techniques (referred to as Kolosov-Muskhelishvili formalism) [20]). Appendix G presents two worked examples, which illustrate the mapping from tension to physical space and the consequence of tension curvature.
A remarkable consequence of these results is that the shape and macroscopic mechanical properties of the tissue are determined by the tension metric and cell density alone. There are many microscopic tension configurations (tension triangulations) that have the same $\mathbf{g}$ and $n$. On a macroscopic level, they are all equivalent – the microscopic details are irrelevant for the static problem. Moreover, up to the overall scale $p_{0}$, the only free parameter of the theory is the cellular bulk modulus $B$.
Let us briefly discuss four special cases: (i) For incompressible cells, $B\rightarrow\infty$ forcing $\partial_{\alpha}u_{\alpha}=-\delta n$. (ii) If pressures follow the ideal gas law $p(n)p_{0}n/n_{0}$, i.e. $B=1$, the total bulk and shear moduli are equal, which may be relevant to experiments on conventional foams with air-filled bubbles. (iii) For $B=1/2$, the total bulk modulus vanishes, so that the conformal mode $f_{\alpha}$ becomes soft. If $B<1/2$, the total bulk modulus is negative: intracellular pressure is insufficient to balance the isotropic tensile stress. Mechanical stability then requires a constraint on the total area. (iv) For fully compressible cells $B=0$, the isogonal potential $\theta$ is an unconstrained soft mode. However, conformal modes, which require pressure gradients, cannot be accommodated. By Eq. (36), freely compressible cells (i.e., $B=0$) are unstable if $K_{g}\neq 0$.
In summary, boundary forces deform the tissue away from the emergent reference state $z(\xi,\bar{\xi})$. Incompatible cell densities and tension curvature lead to residual stresses, present even for free boundaries. The physical stress is governed by effective linear elasticity, leading to the biharmonic Eq. (36) for the isogonal potential, the effective Airy function. Due to Eq. (36), the physical configuration is the one with “minimal stress”, under the constraints imposed by BCs  [^4] Indeed, the biharmonic Eq. (36) can be written variationally: any solution must minimize the functional $\int[(\Delta\theta)^{2}+\theta\,\Delta(B^{-1}\log\lambda_{g}+\delta n)]d^{2}w$. Using $\mathrm{Tr}\bm{\sigma}^{\mathrm{tot}}=\Delta\theta$ and integrating by parts, the stress thus minimizes $\int[(\mathrm{Tr}\bm{\sigma})^{2}+(\mathrm{Tr}\bm{\sigma})(B^{-1}\log\lambda_{g}+\delta n)]d^{2}w$..
.3 Adiabatic dynamics: Beltrami flow
.3.1 Active stress dynamics
In summary, the map $\xi\mapsto w\big(\xi,\bar{\xi}\big)$ embeds the tension metric into physical space and defines the force-balanced state of the tissue. We next consider the temporal dynamics of the physical cell positions $w(\xi,{\bar{\xi}},t)$ in response to time-dependent active tensions. We assume rapid relaxation towards mechanical equilibrium. The tissue thus adopts the instantaneous equilibrium configuration $\mathbf{w}({\bm{\xi}})$, the solution of Eq. (5) (see App. D). We now consider it as a functional $w[\mathbf{g}(t)](\xi,\bar{\xi})$ of the time-dependent tension metric $\mathbf{g}(t)$. In this adiabatic approximation, the tissue flow due to tension dynamics can be computed by the chain rule:
$\partial_{t}w[\mathbf{g}(t)](\xi,{\bar{\xi}})$ $=\frac{\delta w[\mathbf{g}]}{\delta\mathbf{g}}\circ\partial_{t}\mathbf{g}(\xi,\bar{\xi})$
$:=\int\frac{\delta w[\mathbf{g}](\xi,\bar{\xi})}{\delta\mathbf{g}(\xi^{\prime},\bar{\xi}^{\prime})}\,\partial_{t}\mathbf{g}(\xi^{\prime},\bar{\xi}^{\prime})d\xi^{\prime}d\bar{\xi}^{\prime},$ (37)
The functional derivative $\delta w/\delta\mathbf{g}$ is a convolutional integral operator, defined on the second line. This is a consequence of the non-local character of adiabatic flow, where a local change in contractility propagates across the tissue. We evaluate Eq. (37) by the composition of incremental maps and further decompose $w(\xi,\bar{\xi})$ into $\xi\mapsto z(\xi,\bar{\xi})\mapsto w(z,\bar{z})$, from $\xi$ to the reference $z$, and from $z$ to the physical state. By the chain rule:
$\frac{\delta w[\mathbf{g}](\xi,\bar{\xi})}{\delta\mathbf{g}}=\frac{\delta w[\mathbf{g}]}{\delta z}\circ\frac{\delta z[\mathbf{g}]}{\delta\mathbf{g}}+\frac{\delta w[\mathbf{g}]}{\delta\bar{z}}\circ\frac{\delta\bar{z}[\mathbf{g}]}{\delta\mathbf{g}}$ (38)
Here, $\delta z/\delta\mathbf{g}$ describes the variation of the isothermal embedding of $\mathbf{g}$, while $\delta w/\delta z$ involves solving an elasticity problem, under the constraints imposed by the physical BCs (e.g., zero normal stress).
.3.2 Beltrami flow
The dynamics of $z(\xi,\bar{\xi},t)$ is given by the Beltrami Eq. (10). Consider an infinitesimal increment
$\mathbf{g}({\bm{\xi}},t)\mapsto\mathbf{g}({\bm{\xi}},t+dt)=\mathbf{g}({\bm{\xi}},t)+\partial_{t}\mathbf{g}({\bm{\xi}},t)dt$ (39)
The resulting flow $z\mapsto z+\dot{z}\,dt$ must reestablish a state of isotropic stress. Expanding Eq. (10) in $dt$:
$\bar{\partial}_{\xi}\dot{z}(\xi,\bar{\xi},t)=\partial_{t}\mu_{g}(\xi,\bar{\xi},t)$ (40)
subject to the BC $\operatorname{Re}[(\partial_{\xi}z)(\bar{\partial}_{\xi}\dot{z})]|_{\partial\Omega}=\partial_{t}\lambda_{g}|_{\partial\Omega}$. In other words, the anisotropic deformation of the metric $\partial_{t}\mu_{g}$ must be accommodated by a corresponding shear deformation. Tissue flow corresponds to a sequence of infinitesimal QC deformations – a Beltrami flow. The Beltrami flow equation is linear and can be solved by the method of Green’s functions, even for large deformations [26] (Appendix E).
We can restate the flow Eq. (40), formulated in Lagrangian $\xi$-coordinates, in Eulerian $z$-coordinates. Crucially, this allows computing the infinitesimal change in the embedding fully in terms of the present-time reference. We therefore reexpress the flow in the current $z(t)$ coordinates, so that $\dot{z}(z(t),\overline{z(t)},t)$ is the Eulerian velocity field. For brevity, we write $z=z(t)$. Eq. (40) becomes
$\bar{\partial}_{z}\dot{z}(z,\bar{z})=\frac{1}{2\lambda_{g}^{2}}\partial_{t}g_{zz}(z,\bar{z},t)$ (41)
(In the partial time-derivative $\partial_{t}$, the coordinates $z=z(t)$ are kept fixed. By definition of the isothermal $z(t)$, the total time derivative vanishes $\tfrac{d}{dt}g_{zz}(z(t),\bar{z}(t),t)=0$, but $\partial_{t}g_{zz}(z,\bar{z},t)$ does not.)
Eq. (41) can be brought into an alternate form which sheds light on the nature of the adiabatic flow. First expand Eq. (4) around $g_{z\bar{z}}=2\lambda_{g}^{2},\;g_{zz}=0$:
$\sigma_{\bar{z}z}$ $=p_{0}(2\lambda_{g}+(\partial_{t}g_{\bar{z}z})dt)$ (42a)
$\sigma_{zz}$ $=-\frac{p_{0}}{2\lambda_{g}}(\partial_{t}g_{zz})dt=-p_{0}\lambda_{g}(\partial_{t}\mu_{g})dt$ (42b)
(the minus sign $\partial_{t}g_{zz}$ in Eq. (42b) comes from the $\pi/2$-rotation in Eq. (4).) Taking the divergence, Eq. (41) becomes, in Cartesian notation
$\frac{p_{0}}{2}\Delta(\dot{z})_{\beta}=\tfrac{1}{2}\partial_{\beta}(\lambda_{g}^{-1}\partial_{t}\sigma_{\alpha\alpha})-\partial_{\alpha}(\lambda_{g}^{-1}\partial_{t}\sigma_{\alpha\beta}).$ (43)
Crucially, while Eq. (43) resembles a Stokes flow, it is driven not by stress $\sigma$, but by its time-derivative $\partial_{t}\sigma$. The stress increment $(\partial_{t}\sigma_{\alpha\beta})dt$ is in the current configuration $z(t)$, and therefore not necessarily isotropic. In the new $z(t+dt)=z(t)+\dot{z}\,dt$, stress is again isotropic and can be balanced by pressure (an isotropic $\partial_{t}\sigma$ only changes the local pressure and does not induce $\dot{z}$-flow). Moreover, in contrast to a physical Stokes flow, Eq. (43) describes the “flow” of the (stress-free) reference configuration $z$.
.3.3 Physical flow
We are ultimately interested in the dynamics of the physical configuration. The change of the map from reference to physical configuration is $w(z,\bar{z},t+dt)=w(z,\bar{z},t)+\dot{w}(z,\bar{z},t)\,dt$. The physical flow $\dot{w}$ is determined by the BCs and the constitutive relation $P(n)$ (see Sect. .2.4). For instance, for $B=\infty$, one obtains:
Eq. (44a) enforces incompressibility. Eq. (44b) derives from the solvability Eq. (24), expanded for $|w-z|\ll 1,|\lambda_{g}-1|\ll 1$: the physical flow relative to the changing reference configuration must obey the force-balance compatibility condition. For free BCs, Eq. (44b) could be solved simply by $\dot{w}=\dot{z}$. By Eq. (44a), this requires $\dot{z}$ to be incompressible, i.e., $\dot{z}=2i\bar{\partial}_{z}\psi$ for a real-valued solenoidal potential $\psi$. Plugging $\dot{z}=2i\bar{\partial}_{z}\psi$ into Eq. (41), Eq. (13) leads to the condition $\dot{K_{g}}=0$. Hence, for flat changes in $\mathbf{g}$ and free BCs, reference-flow translates directly into physical flow. In general, one must solve the elasticity problem Eq. (44) to determine $w$ in terms of $z$.
Together, Eqs. (40) and (44) describe the reestablishment of force balance after an adiabatic change of stress. Flow establishes force balance, but the tensile stress remains non-zero. In this adiabatic regime, the flow rate is set by the rate of change of the tension metric, i.e. the timescale on which cells regulate contractile force generation. At this point, it would appear that a large shape change requires building up increasingly large tension anisotropy (i.e., anisotropy of tension triangles). As we will see next, however, cell rearrangements plastically lock-in past tension dynamics while resetting the built-up tension anisotropy. Macroscopically, these dynamics appear as a steady active plastic flow driven by an effective active stress.
.4 Cell rearrangement
.4.1 Continuum representation of microscale topology and tension anisotropy
Large tissue deformations cause the cells in a tissue to rearrange, leading to plastic shape changes. In the discrete ATN model, such T1 processes correspond to edge flips that change the adjacency graph of the triangulation (Fig. 4). These flips occur when the length of a cell-cell interface vanishes $\ell_{ij}=0$ and do not manifest themselves in the tension metric. The challenge is to represent the inherently discrete T1 processes in the continuum.
Let us recall that in addition to the tension metric $\mathbf{g}$, the discrete ATN model depends on the topology of the cell tiling, i.e., the vertex adjacency matrix. To capture the adjacency information in the continuum, we introduce an “adjacency metric” $a_{\alpha\beta}(\bm{\xi})$ (defined so that adjacent cells have unit distance). It is the continuum limit of a triangulation with the same adjacency as the tension triangulation, but with all edge lengths set to unity. The metrics $\mathbf{g}(\bm{\xi})$ and $\mathbf{a}(\bm{\xi})$ define two distinct metrics on the same manifold. The Lagrangian coordinates $\bm{\xi}$ are attached to material elements (e.g. cells) and label corresponding points. The adjacency metric “geometrizes” the combinatorial information on cell-cell adjacency, and is closely related to the Koebe-Andreev-Thurston circle packing theorem (see Discussion).
Intuitively, the “discrepancy” between $\mathbf{a}$ and $\mathbf{g}$ measures the shape of tension triangles in the continuum theory, notably, the local tension anisotropy ($\mathbf{a}$ is made from equilateral “reference” triangles, $\mathbf{g}$ from potentially anisotropic tension triangles). Gaussian curvature of $\mathbf{a}$ corresponds to the net density of topological defects in a triangular lattice (i.e., 5- or 7-sided cells, Fig. 4555In an equilateral triangulation, a coordination number $\neq 6$ implies a nonzero angle deficit around a vertex, and hence non-zero Gaussian curvature concentrated at such vertices. The area element of $\mathbf{a}$ defines the local cell density, $n=n_{0}\sqrt{\det\mathbf{a}}$ (we dimensionalize $\mathbf{a}\mapsto\mathbf{a}/n_{0}$ so that $\mathbf{a}$ has units of $\mathrm{length}]^{2}$ [^6] Microscopically, the cell area in an equilateral triangulation is proportional to the coordination number. Hence, strictly speaking, $\mathbf{a}\mapsto\mathbf{a}/n_{0}$ only holds if there is no large-scale coordination number gradient..
Together, the two metrics $\mathbf{a},\,\mathbf{g}$ furnish a complete continuum description of the cell tiling and its mechanics. The ability to control tensions and topology separately is what separates the ATN model from an ordinary fluid foam, where $\mathbf{g}=\mathbf{a}$ (all tensions are equal and can be set to unity).
Cell rearrangements correspond to dynamics in the adjacency metric $\mathbf{a}$. To describe them, we now introduce isothermal coordinates $\zeta(\xi,\bar{\xi})$ for the adjacency metric $\mathbf{a}$, analogous to the $\mathbf{g}$-isothermal coordinates $z$:
$\bar{\partial}_{\xi}\zeta=\mu_{a}\partial_{\xi}\zeta\;\Rightarrow a_{\zeta\zeta}=0,\ \ a_{\zeta\bar{\zeta}}=2\lambda_{a}^{2}$ (45)
The common Lagrangian cell labels $\xi$ establish a one-to-one correspondence between the two isothermal coordinates $z,\;\zeta$ for the tension and adjacency metric:
[…]
Since $\mathbf{a}$ and $\mathbf{g}$ are independent metrics, $\zeta(\xi,\bar{\xi})$ and $z(\xi,\bar{\xi})$ are independent maps. Via Eq. (46), we switch from the $\xi$- to the $\zeta$-basis:
$\xi\xmapsto{\mathbf{a}\text{-iso}}\zeta(\xi,\bar{\xi})\xmapsto{\mathbf{g}\text{-iso}}z(\zeta,\bar{\zeta})\xmapsto{\partial_{\alpha}\sigma_{\alpha\beta}=\partial_{\beta}p}w(z,\bar{z})$ (47)
The adjacency coordinates $\zeta$ serve as a convenient set of coordinates adapted to the adjacency of the cell network – in $\zeta$-coordinates, cells have isotropic shapes and density $n_{0}\lambda_{a}^{2}$. If interfacial tensions are perfectly isotropic, $\mathbf{a}=\mathbf{g}$ and $z=\zeta$. More generally, the Jacobian $\mathbf{D}z(\zeta,\bar{\zeta})$ of the QC map from $\zeta$ to the tension reference state $z$ defines the local tension configuration [29, 16]. Microscopically, $\mathbf{D}z(\zeta,\bar{\zeta})$ deforms equilateral “adjacency” triangles into anisotropic tension triangles. In particular, the Beltrami coefficient $\mu_{\zeta|z}=\bar{\partial}_{\zeta}z/\partial_{\zeta}z=\mu_{g}(\zeta,\bar{\zeta})$ encodes the magnitude and direction of the tension anisotropy. Overall, in $z$-coordinates, the adjacent metric simply specifies the local cell density and tension anisotropy.
By contrast, the map from $z$ to the physical configuration $w$, and its Beltrami coefficient $\mu_{z|w}=\bar{\partial}_{z}w/\partial_{z}w$, determines the elastic shear and macroscopic deviatoric stress. In fact, $\mu_{z|w}$ is proportional to deviatoric stress [cf. Eq. (20)]. The total anisotropy of the map $w(\zeta,\bar{\zeta})$ from isotropic cells to the physical configuration is given by the chain rule:
$\mu_{\zeta|w}$ $=\frac{\bar{\partial}_{\zeta}w}{\partial_{\zeta}w}=\frac{\bar{\partial}_{\zeta}z\partial_{z}w+\bar{\partial}_{\zeta}\bar{z}\bar{\partial}_{z}w}{\partial_{\zeta}z\partial_{z}w+\partial_{\zeta}\bar{z}\bar{\partial}_{z}w}$
$=\frac{\mu_{\zeta|z}+e^{-i2\phi_{\zeta|z}}\mu_{z|w}}{1+e^{-i2\phi_{\zeta|z}}\bar{\mu}_{\zeta|z}\mu_{z|w}}\approx\mu_{\zeta|z}+e^{-i2\phi_{\zeta|z}}\mu_{z|w}$ (48)
where $e^{-i2\phi_{\zeta|z}}=\overline{\partial_{\zeta}z}/\partial_{\zeta}z$, and we assumed $|\mu_{\zeta|z}|,|\mu_{z|w}|\ll 1$ to make the approximation in the last step.
Taken together, the $\mathbf{a}$-isothermal coordinates $\zeta$ allow us to define the local tension configuration in the continuum as the Jacobian of the map $z(\zeta)$. Per Sect. .2, the macroscopic stress is independent of the adjacency metric and hence of tension anisotropy $\mu_{\zeta|z}$. However, the latter plays an important role in cell rearrangement, as we will see below.
.4.2 Kinematics of cell rearrangement
Dynamics of the adjacency metric are driven by the (oriented) T1 rate $R(\xi,\bar{\xi},t)$, to which individual T1s make infinitesimal contributions. These increments are most easily expressed in Eulerian $\zeta(t)$-coordinates
$\mathbf{a}(\bm{\zeta}(t),t)\mapsto\mathbf{a}(\bm{\zeta}(t),t)+\partial_{t}\mathbf{a}(\bm{\zeta}(t),t)dt$ (49)
As above, we write $\zeta=\zeta(t)$ for short in the following. Since T1s do not change the cell number, they correspond to pure shear deformation of $\mathbf{a}$:
$\partial_{t}\lambda_{a}(\zeta,\bar{\zeta},t)=0,\quad\partial_{t}\mu_{a}(\zeta,\bar{\zeta},t)=R(\zeta,\bar{\zeta},t)$ (50)
Eq. (50) can be considered a definition of the rate of topological rearrangement in the continuum.
Figure 4: In an hexagonal lattice, a T1 transition generates two 5-7 defect pairs. Correspondingly, a localized shear generates a curvature quadrupole in the adjacency metric $\mathbf{a}$.
For example, consider a localized “rearrangement event” $da_{\zeta\zeta}=R\,dt=c\ell_{R}^{2}e^{-\zeta\bar{\zeta}/(2\ell_{R}^{2})}\,dt$ with spatial extent $\ell_{R}$, on an initially uniform background $\lambda_{a}=1$. The complex coefficient $c$ sets the shear orientation and magnitude. To mimic an “isolated T1” in a hexagonal lattice, one sets $\ell_{R}$ to the scale of the cell size, and $|c|\sim 1/t_{\mathrm{T1}}$, where $t_{\mathrm{T1}}$ is the timescale of the T1 process. Microscopically, such a T1 creates a 5-7 defect pair (Fig. 4, top). In the continuum, defects are localized quadrupoles in the adjacency curvature $K_{a}$. Indeed, using the St. Venant formula Eq. (13), the increment $da_{\zeta\zeta}$ leads to:
$dK_{a}=\ell_{R}^{-2}\operatorname{Re}\!\left[c\bar{z}^{2}\right]e^{-\zeta\bar{\zeta}/(2\ell_{R}^{2})}\,dt$ (51)
For instance, for $\arg c=0$, $dK_{a}\propto(\zeta_{1}^{2}-\zeta_{2}^{2})\,dt$ (see Fig. 4, bottom).
The T1-induced shear deformation Eq. (50) requires an update to the isothermal $\zeta$-coordinates. The “topological” flow $\zeta(t)\mapsto\zeta(t+dt)=\zeta(t)+\dot{\zeta}(\zeta(t),\bar{\zeta}(t),t)dt$ obeys a Beltrami flow equation:
$\bar{\partial}_{\zeta}{\dot{\zeta}}(\zeta,\bar{\zeta},t)=\partial_{t}\mu_{a}(\zeta,\bar{\zeta},t)=R(\zeta,\bar{\zeta},t)$ (52)
In the updated $\zeta(t+dt)$-coordinates, $\lambda_{a}(\zeta(t+dt),\bar{\zeta}(t+dt))$ changes to accommodate the curvature increment due to $\partial_{t}\mu_{a}(\zeta(t),\bar{\zeta}(t),t)$ in the old $\zeta(t)$ coordinates.
The total time derivative of $\mathbf{g}(\bm{\zeta})$ combines two effects: intrinsic remodeling and the coordinate transformation due to the cell rearrangement flow $\zeta\mapsto\zeta+\dot{\zeta}\,dt$. The transformation law Eq. (1) implies
$\frac{d\mathbf{g}(\zeta,\bar{\zeta})}{dt}=\partial_{t}\mathbf{g}+\left[({\dot{\zeta}}\cdot\mathbf{D})\mathbf{g}-\mathbf{D}{\dot{\zeta}}\cdot\mathbf{g}-\mathbf{g}\cdot(\mathbf{D}{\dot{\zeta}})^{T}\right]$ (53)
The plastic flow of $\zeta$-coordinates appears as a continuous “reparametrization” of the tension manifold. The convective derivative of the Beltrami-coefficient $\mu_{\zeta|z}$, assuming $|\mu_{\zeta|z}|\ll 1$, is given by:
$\frac{d\mu_{\zeta|z}}{dt}$ $\approx\partial_{t}\mu_{\zeta|z}-\partial_{t}\mu_{a}+\operatorname{Re}[\bar{\dot{\zeta}}\partial_{\zeta}]\mu_{\zeta|z}$
$\qquad+2i\operatorname{Im}[\partial_{\zeta}{\dot{\zeta}}]\,\mu_{\zeta|z}$ (54)
The first two terms describe the composition of two QC transformations: by (52), $\partial_{t}\mu_{a}$ is the Beltrami coefficient of the infinitesimal map $\zeta(t)\mapsto\zeta(t+dt)$ (see Appendix F.1 for details, and Fig. 5 for an illustration of the underlying geometry). The last two terms correspond to advection and co-rotation.
To summarize the kinematic logic: T1s change the adjacency metric $\mathbf{a}$, necessitating an update of the isothermal coordinates $\zeta$. This corresponds to a reparametrization of the tension metric $\mathbf{g}$, i.e. a change of $\mu_{\zeta|z}$. This coordinate transformation leaves the tension manifold (i.e., $\mathbf{g}({\bm{\xi}})$) invariant. In our framework, cell rearrangement, thus, is a purely kinematic process and has no instantaneous effect on $\mathbf{g}$. This is best motivated by looking at physical space. Microscopically, a T1 flips an edge whose physical length $\ell_{ij}=0$, and is hence “invisible”: it has no instantaneous effect on physical shape or macroscopic stress. A specific model for the dynamic change of the tension manifold elicited by the kinematic effect of T1s on tension anisotropy will be described in the next section.
.4.3 Dynamics of cell rearrangement: Active and passive T1s
To fill the kinematic relations (52)–(.4.2) with life, we need to specify a continuum law for the rate and orientation of T1s $\partial_{t}\mu_{a}$ and a model for the dynamics of $\mathbf{g}$.
As a minimal, phenomenological model for topological plasticity, we propose to represent the effect of T1s by the continuous relaxation of the total anisotropy $\mu_{\zeta|w}$. The (oriented) rearrangement rate, therefore, reads
$\partial_{t}\mu_{a}(\zeta,\bar{\zeta})$ $=-\gamma_{\mathrm{T1}}\frac{\delta|\mu_{\zeta|w}|^{2}}{\delta\bar{\mu}_{a}}=\gamma_{\mathrm{T1}}\mu_{w|z}$
$\approx\gamma_{\mathrm{T1}}(\mu_{\zeta|z}+e^{-i2\phi_{\zeta|z}}\mu_{z|w})$ (55)
Here, $\gamma_{\mathrm{T1}}$ is a rate coefficient (the “T1 rate”), which depends on the distribution of microscopic tension configurations, as we discuss below. The variation of $|\mu_{\zeta|w}|^{2}$ follows from the Beltrami composition formula (48) (see Appendix F.2). Equation (.4.3) implies that T1s can be driven by elastic deformations due to external forces acting on the tissue boundary, or by internal changes of tensions [29, 16]. We refer to these two scenarios as passive and active T1s. Passive T1s are driven by anisotropic elastic strain $\mu_{z|w}$ (and hence stress anisotropy $\sigma$). By contrast, active T1s can occur in the absence of macroscopic stress. Instead, they are driven by the tension anisotropy $\mu_{\zeta|z}$.
Thus, Eq. (.4.3) qualitatively reproduces the mechanisms driving T1s as revealed by geometric considerations in the discrete setting [29, 15]. There, the combined tension (active) and strain (passive) anisotropy determines when a cell-cell interface reaches zero length $\ell_{ij}=0$, triggering a T1. The critical microscopic Beltrami coefficient for which $\ell_{ij}=0$ can be worked out geometrically as $|\mu_{\zeta|w}^{\triangle}|=1/2$ [15]. Here, the superscript $\triangle$ marks quantities on the single-triangle level. The Beltrami coefficients $\mu_{\zeta|z}$ and $\mu_{z|w}$ entering the continuum theory represent a macroscopic average over many tension triangles. The microscopic T1 threshold $|\mu_{\zeta|w}^{\triangle}|=1/2$ implies the existence of a yield strain for passive T1s: below a critical $|\mu_{z|w}|$, the T1 rate $\gamma_{\mathrm{T1}}$ vanishes. More generally, $\gamma_{\mathrm{T1}}$ depends on the microscopic distribution of cell-level configurations $\mu_{\zeta|w}^{\triangle}$, which must be summarized by a suitable set of local order parameters. A mean-field approach [30] then allows systematic calculation of $\gamma_{\mathrm{T1}}$ in terms of these order parameters, together with equations governing their dynamics.
.4.4 T1s relax tension and stress anisotropy
To complete the model for plasticity in the continuum, we need to address how cell rearrangements impact the tension metric $\mathbf{g}$. A priori, the change of the adjacency metric $\mathbf{a}$ corresponds to a reparametrization of $\mathbf{g}$ – the tension manifold remains invariant under cell rearrangements (the macroscopic stress depends only on $\mathbf{g}$ and the physical BCs, also remains invariant). Edge flips occur in the triangulation pass between different triangulations that belong to the same stress state. However, T1s are not only an abstract “remeshing” of the tension manifold – they create a new cell interface whose tension dynamics are controlled by specific (bio-)physical processes at the microscopic level.
More precisely, the edge flips kinematically change the tension anisotropy. This is geometrically evident in the discrete setting (as illustrated in Fig. 5 in App. F). In the continuum, where tension anisotropy is measured by $\mu_{\zeta|z}$, the chain rule leads to a contribution $\tfrac{d}{dt}\mu_{\zeta|z}\sim-\partial_{t}\mu_{a}$ in Eq. (.4.2). The T1-induced change of tension anisotropy elicits subsequent dynamics of $\mathbf{g}$ through the processes that govern the microscopic tensions.
Consider the case of a conventional fluid foam as the paradigmatic case for passive T1s. External forces build up macroscopic stress $\sigma$ through isogonal strain while microscopic tensions remain constant. The isogonal strain will cause T1s, which kinematically build up tension anisotropy that cannot be sustained since surface tension is fixed. Instead, the fluid in the interfaces rapidly redistributes, leading to transient viscous stresses and relaxation back to $\mu_{\zeta|z}=0$ concomitant with a relaxation of macroscopic stress $\sigma$. Passive T1s therefore reproduce Maxwell stress relaxation. In the limit where the microscopic timescale on which viscous stresses relax is zero, $\mathbf{g}\equiv\mathbf{a}$ and $\zeta=z$ in a conventional foam. Therefore,
$\partial_{t}\mu_{g}=\partial_{t}\mu_{a}=\gamma_{\mathrm{T1}}\mu_{z|w}\approx\gamma_{\mathrm{T1}}\frac{\sigma_{ww}}{2p_{0}}$ (56)
since deviatoric strain $\mu_{z|w}$ determines the stress via Eq. (.2.4) (of the the linearized theory). On the other hand, tension metric dynamics determines the change in stress. Indeed, for fixed physical positions $w$, the Beltrami equation (10) implies $\frac{\delta\mu_{z|w}}{\delta\mu_{g}}=-1$, and therefore, $\partial_{t}\sigma_{ww}=-\partial_{t}\mu_{g}/(2p_{0})$. Together, these two effects result in effective Maxwell relaxation $\partial_{t}\sigma\sim-\sigma$. (While T1s relax only the anisotropic stress, due to mechanical balance Eq. (22), this also removes isotropic stress gradients.) For the reference-state velocity field in the purely passive case, stress relaxation and the “adiabatic flow” Eq. (43) lead to an effective Stokes flow:
$\frac{p_{0}}{2}\Delta\dot{z}_{\beta}=\gamma_{\mathrm{T1}}\partial_{\alpha}\!\left(\sigma^{\mathrm{tot}}_{\alpha\beta}-\tfrac{1}{2}\sigma^{\mathrm{tot}}_{\gamma\gamma}\delta_{\alpha\beta}\right)$ (57)
The drive results from the anisotropic macroscopic stress, and the T1 rate sets the relaxation rate. Going from $\dot{z}_{\beta}$ to the physical flow $\dot{w}_{\beta}$ requires additionally solving Eq. (44). If $\dot{K}_{g}=0$, one simply has $\dot{w}_{\beta}=\dot{z}_{\beta}$.
Let us next consider the purely active case, $w=z$, i.e. $\mu_{z|w}=0$. Here, T1s happen when $|\mu_{\zeta|z}|$ increases either by positive feedback [29, 16] or by an “external field” like a morphogen gradient [31, 30]. Substituting Eq. (.4.3) into Eq. (.4.2) reveals that T1s kinematically relax tension anisotropy: $\tfrac{d}{dt}\mu_{\zeta|z}=\partial_{t}\mu_{\zeta|z}-\gamma_{\mathrm{T1}}\mu_{\zeta|z}$ (suppressing the advection and co-rotation terms). Steady active plastic flow results when driving and relaxation balance at a finite T1 rate. (Appendix G.3 describes a concrete model where tensions are governed by positive mechanical feedback [30].) Active flow can take place at vanishing macroscopic stress and is therefore fundamentally different from conventional Maxwell viscoelasticity. Instead, it is characterized by sustained microscopic tension anisotropy, experimentally detectable as anisotropic junctional myosin [32].
We conclude that both active and passive T1s relax their “driving anisotropy”: tension anisotropy in the former case and macroscopic stress anisotropy in the latter. We can combine both in a unified phenomenological model for the T1-contribution to the dynamics of $\mathbf{g}$
$\left(\partial_{t}\mu_{g}(\zeta,\bar{\zeta})\right)_{\mathrm{T1}}=\gamma_{\mathrm{T1}}e^{-i2\phi_{\zeta|z}}\mu_{z|w}=-\gamma_{\mathrm{T1}}\frac{\delta|\mu_{z|w}|^{2}}{\delta\bar{\mu}_{g}},$ (58)
Where the second equality follows from Eq. (.4.3). Additional contributions can arise from mechanical feedback and “external fields” driving $\mu_{g}$, while in a passive tissue (or conventional foam), only T1s contribute.
To summarize, the present framework describes a tissue by two metrics $\mathbf{a},\mathbf{g}$ that can be controlled independently. A kinematical prescription determines what happens during topological change (dynamics of $\mathbf{a}(t)$, keeping $\mathbf{g}(t)$ fixed). The dynamics of the tension metric $\mathbf{g}$ depend on the specific model or biological context of interest. We showed that minimal and microscopically motivated choice, Eq. (58), naturally gives rise to Maxwell-like stress relaxation through passive T1s while active T1s relax tension anisotropy.
Discussion
We have presented a continuum theory for active tension nets, a paradigmatic example of “active elasticity”, where the local stress, rather than the local target shape, is the physical (or biological) control parameter. We showed how the intrinsic configuration of active stress controls the large-scale shape of a tissue. Adiabatic changes of stress drive active plastic flow, described as a QC flow governed by two coupled Beltrami equations, Eqs. (40) and (52).
From cell-level tensions to macroscopic stress.
We have shown that the microscopic configuration of active tensions $\tau_{ij}$ defines, in the continuum limit, a Riemannian tension metric $\mathbf{g}$. This metric, together with the mechanical BCs, determines the macroscopic stress tensor $\sigma$, and the positions $w$ adopted by cells in mechanical equilibrium. By identifying the tension metric $\mathbf{g}$ as the relevant macroscopic degree of freedom of active tension nets, our theory explains how local contractile activity translates into large-scale shape.
Crucially, many microscopic tension triangulations specify the same tension metric. They correspond to reparametrizations of the tension manifold and do not affect physical stress. Put differently, the macroscopic (i.e., tissue-scale) mechanical state only depends on the shape of the tension manifold, not how it is discretized by “painting” individual tension triangles onto it, like in a finite-element discretization of a partial differential equation. In particular, different triangulations of a region in the plane all define the same tension metric – only the shape of the region’s boundary is relevant. In that sense, for a flat tension metric, the continuum theory is “holographic”: the configuration of the tissue is encoded in the shape of the boundary and the traction forces acting there. This holographic nature of emergent elasticity is also relevant in granular matter, where boundary forces constrain the ensemble of self-stresses in the bulk [33].
Of course, the microscopic cell shapes depend on the tension triangulation. Moreover, tensions are ultimately sensed and controlled at the cellular level. Therefore, their dynamics depend on the microscopic tension configuration. A continuum model, therefore, requires information in addition to $\mathbf{g}$. The most important piece is the adjacency metric $\mathbf{a}$, whose continuum dynamics captures cell rearrangements. While the tension metric $\mathbf{g}$ determines the macroscopic mechanics, the local tension anisotropy is encoded in the “difference” of $\mathbf{g}$ and $\mathbf{a}$ – the shape of tension triangles. Overall, the present theory comprises the minimal description of tension nets in the continuum: The tension metric $\mathbf{g}({\bm{\xi}})$ (equivalently, isothermal embedding $z(\xi,\bar{\xi})$ and curvature $K_{g}$) and the adjacent metric $\mathbf{a}({\bm{\xi}})$, for a total of six degrees of freedom. The systematic description of stress and network topology in terms of Riemannian geometry and QC maps may prove useful in the study of other 2D materials with dynamic microscopic structure, such as granular matter.
Active plastic flow.
To describe tissue flow driven by adiabatic dynamics of active stress, one tracks the reference state $z(\xi,\bar{\xi},t)$ of the evolving tension metric $\mathbf{g}(t)$. The time evolution $\dot{z}$ is governed by a Beltrami flow, Eq. (40). Large deformations entail topological cell rearrangement. This raises the challenge of incorporating discrete topology into a continuum theory. On a cell level, topological information can be “geometrized” via the so-called circle packing of the cell adjacency graph, whose existence and uniqueness is guaranteed by the Koebe–Andreev–Thurston theorem [34]. As explained in the companion paper [15], the circle packing is a “discrete” conformal embedding of the equilateral triangulation formed by the adjacency graph. In the continuum, it becomes the isothermal embedding $\zeta(\xi,\bar{\xi})$ of the adjacency metric $\mathbf{a}$. Tension dynamics, plastic flow, and elastic deformation can thus be formulated as two coupled Beltrami flows, Eqs. (40), and (52). This principled approach naturally incorporates the local tension configuration via the Jacobian $\partial_{\bm{\zeta}}\mathbf{z}$.
In conventional models of tissue flow, imbalanced active stresses act against viscous dissipation [32, 35, 36]. In the present theory, by contrast, it is the adiabatic change of active stress that drives flow. This has implications for connecting cell-level models to their large-scale behavior. In this picture, a steady-state tension anisotropy emerges from the balance of anisotropy-increasing processes (such as mechanical feedback or morphogen “drive”) and relaxation by T1s. By contrast, the macroscopic stress remains isotropic (in the absence of external forces). Tension anisotropy has previously been identified with an anisotropic active stress tensor [32]. These discrepant predictions could be tested by laser ablations on a supracellular scale.
In the “active plasticity” model, the tissue remodels from one solid configuration to the next – it flows while remaining solid, able to resist external shear forces [^7] Note, however, that the yield strain becomes highly anisotropic during active plastic flow. Due to the ongoing T1 transitions, there is a large number of cells near the T1 threshold, and the yield strain parallel to tension anisotropy can become very small.. By contrast, in a conventional active fluid, flow results from viscous dissipation, and the system is instantaneously fluid and cannot resist shear forces.
Emergent metric elasticity.
On a formal level, the present theory has parallels with metric elasticity, where a Riemannian reference metric represents the rest lengths between adjacent material points [38], and deviations from these rest lengths in the physical configuration determine the elastic stress via a constitutive law. By contrast, the tension metric specifies the active stress (tension flux) between adjacent points. A stress-free reference state and an effective stress-strain relationship then emerge from the intrinsic stress configuration. The physical stress tensor is determined by the embedding of the tension metric given by the cell positions, i.e., the arrangement of active force dipoles in physical space. The present theory, therefore, furnishes the continuum limit of cell-level models dominated by active mechanics. By contrast, conventional metric elasticity is well-suited to describe the continuum limit of passive models (generalized spring-networks). For instance, in the context of the well-known “area-perimeter” vertex model [7], Ref. [8] defined a network tensor that plays the role of the elastic reference metric. Plastic flow can be modeled by relaxational dynamics of the reference metric [39].
An attractive feature of metric elasticity is its variational formulation. Our theory also admits a variational form: The isothermal embedding $\mathbf{z}({\bm{\xi}})$ of the tension metric can be characterized by a pseudo-energy, formally similar to metric elasticity. Since $\mathbf{z}({\bm{\xi}})$ is a conformal map from the tension manifold into the Euclidean plane, its components must be harmonic functions. Equivalently, $z(\xi,\bar{\xi})$ must be a minimum of the Dirichlet pseudo-energy $E_{D}=\int d^{2}\xi\sqrt{g}\ g^{\alpha\beta}\partial_{\alpha}{\bar{z}}(\xi)\partial_{\beta}{z}(\xi)$ [40] (with suitable BCs, see App. C). However, $E_{D}$ derives from the geometry of force balance, not from a material law, and does not determine the stress tensor $\sigma_{\alpha\beta}$. A better understanding of the variational formulation of active mechanics is an interesting avenue for future work.
Tissues on curved and dynamic surfaces.
For simplicity, the present manuscript considers a tissue embedded in the plane. However, the geometric nature of the theory implies that it easily generalizes to tissues on curved surfaces, such as the ellipsoidal blastoderm of the Drosophila. To this end, the Euclidean metric $\delta_{\alpha\beta}$ in Eq. (11) must be replaced by the metric of the curved physical space. This is most conveniently done by parameterizing physical space with (yet another set of) isothermal coordinates. The reference configuration $z(\xi,\bar{\xi})$ is found as a conformal map between two curved surfaces – the tension- and physical manifold – and the mechanics is controlled by their relative geometry – e.g., the difference of Gaussian curvatures in Eq. (12).
In the above setting, the cells are still constrained to lie on a fixed (albeit curved) physical surface. More generally, one can consider tissues that form dynamic surfaces, embedded in three-dimensional space. One must additionally account for normal force balance, which couples in-plane stress and extrinsic physical curvature to normal pressure gradients [3]. This generalization is required in situations where in-plane contractility drives three-dimensional shape change of tissue sheets, e.g., during visceral organ morphogenesis [41].
Pressure effects and mechanical feedback.
Gaussian curvature of the tension metric leads to pressure differentials, similar to residual stresses in metric elasticity due to an incompatible target metric. Interestingly, many tissues can avoid large pressure differentials and the potential instabilities they entail, like cell extrusion or rupture. On the cell level, pressures can be locally controlled via (negative) feedback loops  [^8] The pressure feedback discussed here operates quasi-statically and requires a non-zero bulk modulus to build up pressure gradients. By contrast, in a compressible tissue, tension incompatibility causes secular elongation or contraction of cell-cell interfaces, requiring a different stabilizing mechanism (see Ref. [11]). coupling tensions and intracellular pressure. A minimal mathematical model for such feedback is $\partial_{t}\mathbf{g}=-\beta(\Delta\log p/p_{0})\mathbf{g}=-\beta K_{g}\,\mathbf{g}$, where $\beta$ is a feedback coefficient and we used Eq. (12). Cells can hence implement a “flattening” Ricci flow [43] by modulating their overall contractile activity in response to intracellular pressure. In contrast to, e.g., Frank elasticity in liquid crystals, Ricci flow only acts on the “curvature part” of $\mathbf{g}$ – the tension manifold remains free to remodel in-plane to drive morphogenetic flows.
On the tissue level, pressure can be controlled via feedback through cell proliferation and growth [44, 1, 45], which in our formulation would act on the target cell density $n({\bm{\xi}})$ (Eq. (17)). Analyzing such “adaptive” feedback loops, and the combination of robustness and plasticity they confer, is an interesting route for future research.
More generally, our findings have interesting consequences for how cells can sense and react to mechanical cues. We showed that the macroscopic stress $\sigma$ is independent of the local tension configuration, and, therefore, “invisible” on a cell level. Instead, the cell-level “observables” are the tension anisotropy $\mu_{\zeta|z}$ and the intracellular pressure $p$. Cell shape anisotropy could be sensed by intracellular structures such as the nucleus, intermediate filaments, or microtubules.
Link to the microscopic ATN model.
The link between the continuum theory and cell level (the discrete ATN model) is discussed in detail in a companion paper [15]. There, we exploit a theory of discrete conformal maps [46, 47] to bridge the gap between the continuum theory of the present manuscript and the mechanics at the cellular level. This theory constructs a discrete equivalent of the isothermal embedding $z(\xi,\bar{\xi})$. The isogonal (curl-free) and conformal modes identified by the continuum analysis reappear as the geometric soft modes which respect the microscopic force balance constraints [48, 11, 49], while the Liouville equation for the curvature of the tension manifold leads to a generalized von Neumann law for the cell pressures. Coarse-graining the microscopic stress confirms the stress-transformation law Eq. (21). Thus, the “top-down” analysis presented here is independently confirmed by “bottom-up” coarse-graining.
Similarly, (active) plasticity, mediated by cell rearrangements, is captured by remodeling of the cell-adjacency graph. In the continuum, this graph is coarse-grained to the adjacency metric $\mathbf{a}$. The circle packing provides a discrete equivalent to the $\mathbf{a}$-isothermal coordinates (see paragraph on “active plastic flow” above). In the discrete setting, the condition for T1 transitions can be calculated explicitly from purely geometric considerations (a generalized “Delaunay criterion”), justifying the phenomenological T1 rate $\propto\mu_{\zeta|w}$ in amorphous tissue with negligible micro-structural order (Eq. (58)). Micro-structural order can strongly influence T1 transitions, in which case additional order parameters become relevant [16]. A continuum theory for these order parameters and their relation to topological defects in the microscopic topology is an interesting direction for future research.
Outlook.
While we have introduced our theory in the context of epithelial tissue mechanics, we believe it applies more broadly. Non-epithelial tissues, like those composed of protrusive, mesenchymal cells, do not form tension nets in the strict ATN sense, but their architecture and shape are also controlled by balancing active contractile stresses against bulk elasticity. Beyond the biological setting, 2D fluid or ferro-fluid foams [50, 13] correspond to the limiting case where all $\tau_{ij}=\tau=\mathrm{const}$. Tension dynamics could be introduced, for instance, through chemical reactions within foam bubbles. As explained in the companion paper [15], there is also a close connection to granular materials and truss networks, which can be understood as the mechanical Legendre dual of tension nets. In granular materials, one typically studies the ensemble of compatible (force-balanced) stress configurations for a given physical arrangement of particles. This is, in a sense, the dual of the problem considered in the present study. In future work, the Legendre duality might, therefore, provide a starting point for building a continuum theory for granular matter. Indeed, previous phenomenological approaches postulate that granular materials can also be described by “emergent elasticity” [51].
To conclude, we believe that our theory provides a general framework for the emergent elasticity and plasticity of matter whose microscopic mechanics is not governed by conventional constitutive laws but by independently controlled active forces. It quantitatively explains how these local active forces generate large-scale shape.
Acknowledgements.
We thank D. Cislo, A. Košmrlj, and H. Weyer for valuable feedback on the manuscript. B.I.S. acknowledges support of the NSF Physics (PoLS) grant #2210612. N.H.C. is supported by a PCTS fellowship. F.B. acknowledges funding from the Max Planck Society and the Gordon and Betty Moore Foundation post-doctoral fellowship (grant #2919).
Appendix A Change of curvature radius by a conformal map
Here, we compute the change of a line’s curvature radius $R$ by a conformal map $z\mapsto f(z)$ [40]. Consider a line $z(t)$ with normal $\mathbf{n}$}. The rotation of the local normal is given by the vorticity $\omega=\arg\partial_{z}f=\partial_{z}f/|\partial_{z}f|$. Moving a step $dt$ along the transformed curve $f(z(t))$, the tangent $i\mathbf{n}$ rotates by an angle $d\phi=\omega(f(z(t+dt))-\omega(f(z(t))=(i{\mathbf{n}})\cdot 2\bar{\partial}_{z}\omega dt$. Due to the scaling factor $|\partial_{z}f|$, the arc length is $|\partial_{z}f|dt$. Putting this together, the curvature radius reads:
$\frac{1}{R}$ $=\frac{d\phi}{d(\lambda t)}=\frac{1}{|\partial_{z}f|}(i{\mathbf{n}})\cdot 2\bar{\partial}_{z}\arg\partial_{z}f$
$={\mathbf{n}}\cdot 2\bar{\partial}_{z}|\partial_{z}f|$ (59)
Appendix B The isogonal-conformal decomposition
Here, we show that any vector field $u_{\alpha}$ that fulfills the solvability condition $\Delta(\epsilon_{\alpha\beta}\partial_{\beta}u_{\alpha})=0$ (the vorticity is harmonic) can be decomposed as $u_{\alpha}=\partial_{\alpha}\theta+f_{\alpha}$, where $f_{\alpha}$ is a conformal vector field. In contrast to the familiar Helmholtz decomposition into curl- and divergence-free parts, this decomposition is only possible if the solvability condition is fulfilled. For example, a solenoidal vector field $u_{\alpha}=\epsilon_{\alpha\beta}\partial_{\beta}\psi$ must have $\Delta^{2}\psi=0$.
We first aim to identify $\theta^{(1)}$ so that $f_{\alpha}^{(1)}=u_{\alpha}-\partial_{\alpha}\theta^{(1)}$ is harmonic. We must solve $\Delta\partial_{\alpha}\theta^{(1)}=\Delta u_{\alpha}$. This is solvable if the curl $\epsilon_{\alpha\beta}\partial_{\beta}$ of the RHS vanishes, which is the condition we assumed. Next, we aim to make the two components $f_{1},f_{2}$ not just harmonic, but conjugate harmonic to fulfill the Cauchy–Riemann equations. We shift again $f^{(1)}_{\alpha}\mapsto f^{(2)}=f^{(1)}_{\alpha}-\partial_{\alpha}\theta^{(2)}$. The Cauchy–Riemann equations imply that the deviatoric strain of $f^{(2)}$ must vanish. Hence, $\theta^{(2)}$ must fulfill
$\partial_{\alpha}\partial_{\beta}\theta^{(2)}=\frac{1}{2}\left(\partial_{\alpha}f^{(1)}_{\beta}+\partial_{\beta}f^{(1)}_{\alpha}-\partial_{\gamma}f^{(1)}_{\gamma}\delta_{\alpha\beta}\right)$ (60)
Note that the curl of the LHS is $\epsilon_{bc}\partial_{\gamma}\partial_{\alpha}\partial_{\beta}\theta^{(2)}=0$. For Eq. (60) to be solvable, the curl of the RHS must also vanish:
$\epsilon_{\beta\gamma}\partial_{\gamma}\left(\partial_{\alpha}f^{(1)}_{\beta}+\partial_{\beta}f^{(1)}_{\alpha}-\partial_{\delta}f^{(1)}_{\delta}\delta_{\alpha\beta}\right)$ (61)
$=(\epsilon_{\alpha\gamma}\partial_{\gamma}\partial_{\beta}-\epsilon_{\beta\gamma}\partial_{\gamma}\partial_{\alpha})f^{(1)}_{\beta}=\epsilon_{\alpha\beta}\partial_{\gamma}^{2}f_{\beta}=0$
since $f^{(1)}_{\beta}$ is already harmonic. Hence, we arrive at the desired conformal-isogonal decomposition $u_{\alpha}=\partial_{\alpha}\theta+f^{(2)}_{\alpha}=\partial_{\alpha}(\theta^{(1)}+\theta^{(2)})+(u_{\alpha}-\partial_{\alpha}\theta^{(1)}-\partial_{\alpha}\theta^{(2)})$.
Appendix C Boundary conditions
Here, we show that the two modes compatible with force balance, the curl-free and conformal modes, are sufficient to accommodate arbitrary mechanical BCs. Indeed, the curl-free mode is already sufficient to accommodate arbitrary boundary displacements $\mathbf{u}_{\partial\Omega}$ (except for global rotations, which are conformal). Write $\mathbf{u}_{\partial\Omega}(s)=u_{s}\hat{\mathbf{s}}+u_{n}\hat{\mathbf{n}}$, where $s$ is the arc length and $\hat{\mathbf{s}},\hat{\mathbf{n}}$ are the tangent and normal. Since $\mathbf{u}=\nabla\theta$ the isogonal potential must fulfill
$\partial_{s}\theta=u_{s},\quad\partial_{n}\theta=u_{n}$ (62)
along the boundary. We can therefore find $\theta$ along the boundary by integrating $u_{s}$
$\theta|_{\partial\Omega}(s)=\int_{0}^{s}\mathrm{d}s^{\prime}\,u_{s}(s^{\prime})$ (63)
Via a global rotation, we can ensure the requirement that
$\oint\mathrm{d}s\,u_{s}=\oint\mathrm{d}\mathbf{s}\cdot\mathbf{u}^{(b)}(s)=0$ (64)
To fulfill $\partial_{n}\theta=u_{n}$ we continue $\theta$ into the bulk with slope $u_{n}$, i.e.
$\theta(\mathbf{b}(s)+dn\,\mathbf{n})=\theta|_{\partial\Omega}(s)+dn\,u_{n}(s),$ (65)
where $\mathbf{b}(s)$ parametrizes the boundary curve and $dn$ is the (infinitesimal) distance from the boundary.
By contrast, conformal deformations alone are insufficient. A flat conformal map $f$ is fully specified by its conformal factor $|\partial_{z}f|$, which obeys the Liouville equation $\Delta\log|\partial_{z}f|=0$. This Poisson problem admits prescribing either Dirichlet conditions $|\partial_{z}f||_{\partial\Omega}$, or Neumann conditions $\partial_{\mathbf{n}}|\partial_{z}f||_{\partial\Omega}$ (the boundary curvature ), but not both simultaneously. Equivalently, one can specify BCs for one of the two components $f_{\alpha}$ – the second component is fixed by the Cauchy-Riemann equations $\partial_{\alpha}f_{1}=\epsilon_{\alpha\beta}\partial_{\beta}f_{2}$.
As an example, consider the unit disk in complex polar coordinates $z=re^{i\phi}$. Conformal maps are holomorphic functions $f$ and, hence, analytic, $f(z)=\sum_{n\geq 0}\tfrac{f^{(n)}}{n!}z^{n}$. Importantly, the sum is restricted to $n\geq 0$: hence, the boundary value $f(e^{i\phi})$ has only non-negative Fourier coefficients. A boundary displacement with negative Fourier coefficients cannot be extended to a conformal deformation of the disk. For example, pure shear $\mathbf{u}|_{\partial\Omega}\propto e^{-i\phi}=x\hat{\mathbf{x}}-y\hat{\mathbf{y}}$ is incompatible with a conformal deformation.
Boundary conditions for the Beltrami equation
In the context of the Beltrami equation Eq. (10), the BCs fix the conformal contribution to $z(\xi,\bar{\xi})$. Therefore, the Beltrami equation also admits a single BC only, e.g., the natural BC for $|\partial_{\xi}z|$, Eq. (14). Indeed, the example Appendix G.1 makes it clear that it is impossible to make the tension metric isotropic while keeping the boundaries clamped. The Beltrami equation must be contrasted with the elasticity problem for the physical configuration $w$, which admits two boundary conditions, as we showed above.
Appendix D Adiabatic approximation
In the main text, we treat tissue dynamics within the adiabatic approximation. Consider relaxation dynamics of the cell positions $\mathbf{r}$
$\gamma^{-1}\partial_{t}r_{\alpha}(\xi,{\bar{\xi}},t)=\partial_{\alpha}\sigma_{\alpha\beta}(t)-\partial_{\beta}p(t)$ (66)
where $\gamma$ is a friction coefficient. The stress $\sigma$ depends both on $\mathbf{r}(t)$ and $\mathbf{g}(t)$ by Eq. (4), and is thus time-dependent. The timescale of relaxation to equilibrium is given by $t_{\gamma}=L\gamma/p_{0}$, where $L$ is the system length scale. In the limit where $t_{\gamma}$ is much faster than any other timescale (notably, that of tension dynamics $\partial_{t}\mathbf{g}$), Eq. (66) reduces to force balance:
$\partial_{\alpha}\sigma_{\alpha\beta}(t)-\partial_{\beta}p(t)=0$ (67)
The cells, therefore, adopt the instantaneous equilibrium configuration $\mathbf{w}({\bm{\xi}},t)$, whose time-dependence results from the dynamics of the “parameter” $\mathbf{g}(t)$.
Appendix E Integrating the Beltrami flow via Green’s functions
The Beltrami flow Eq. (40) describes the flow $v(z,\bar{z})$ of the stress-free reference state $z$ due to the dynamics of the tension metric. Note that for any solution $v(z,\bar{z})$ of Eq. (40), $v(z,\bar{z})+f(z)$ for a conformal $f(z)$ is also a solution. The BCs fix this conformal component. In the main text, we used natural BCs, $\lambda_{g}|_{\partial\Omega}=1$. This minimizes the deformation of the embedding $z(\zeta,\bar{\zeta})$ and the pressure gradients in the reference state, providing a convenient starting point for linearization (Eq. (.2.4)). The shape of the natural domain boundary $\partial\Omega$ is, however, time-dependent.
Instead, one can fix the conformal component by demanding that the isothermal coordinates map the tension manifold to the unit disk. We denote these “disk” isothermal coordinates by $\hat{z}$, with $D=\{|\hat{z}|\leq 1\}$. Then, Eq. (40) can be solved by a fixed Green’s function [26]:
[…]
$G(\hat{z},\hat{z}^{\prime})$ $=\frac{\hat{z}(\hat{z}-1)}{(\hat{z}^{\prime}-\hat{z})\hat{z}^{\prime}(\hat{z}^{\prime}-1)}$ (69)
Eq. (68) also allows the explicit computation of the topological flow due to local cell rearrangement. In summary, the dynamics of the reference state becomes a QC flow $v_{(\hat{z})}$ on the disk $D$, and is analytically tractable. The price to be paid is that the deformation from the reference to the physical state is now, in general, very large.
[…]
Denoting the conformal factor by $\lambda_{\hat{z}|z}=|\partial_{\hat{z}}z|$, the natural BCs require $\lambda_{\hat{z}|z}^{2}|_{\partial_{D}}=\lambda_{g}^{2}(\hat{z},\bar{\hat{z}})|_{\partial_{D}}$ on the disk boundary. We now note that $\Delta\log\lambda_{\hat{z}|z}=0$, since $\log\lambda_{\hat{z}|z}$ is the real part of the analytic function $\log(\partial_{\hat{z}}z)$. Therefore, we can obtain $\log\lambda_{\hat{z}|z}$ on the entire disk by convolution with the Poisson kernel $P(\hat{z},\hat{z}^{\prime})$ as
[…]
$P(\hat{z},\hat{z}^{\prime})$ $=\operatorname{Re}\left[(\hat{z}+\hat{z}^{\prime})/(\hat{z}-\hat{z}^{\prime})\right]$ (71)
In particular, the normal derivative on the boundary is $\partial_{\mathbf{n}}\log\lambda_{\hat{z}|z}|_{\partial D}=(2\pi)^{-1}\int_{\partial D}d\hat{z}^{\prime}\log\lambda_{\hat{z}|z}(\hat{z}^{\prime})\operatorname{Re}\left[\frac{\hat{z}\hat{z}^{\prime}}{(\hat{z}-\hat{z}^{\prime})^{2}}\right]$. Together, $\lambda_{\hat{z}|z}|_{\partial D}$ and $\partial_{\mathbf{n}}\lambda_{\hat{z}|z}|_{\partial D}$ fully describe the shape of the boundary: $\lambda_{\hat{z}|z}$ is the local arc length, and $\kappa=\lambda_{\hat{z}|z}^{-1}+\partial_{\mathbf{n}}\lambda_{\hat{z}|z}^{-1}$ is the curvature (Appendix A). The orientation of the tangent to the boundary at arc length $s$ is given by $\phi(s)=\int_{0}^{s}\kappa(s^{\prime})ds^{\prime}$. One can then integrate to obtain the boundary curve $z(\hat{z})|_{\partial D}$. Finally, the value in the interior of the disk is determined by Cauchy’s integral:
$z(\hat{z})=\frac{1}{2\pi i}\oint_{\partial D}\frac{z(\hat{z})}{\hat{z}^{\prime}-\hat{z}}d\hat{z}^{\prime}.$ (72)
Appendix F Kinematics of Beltrami flow
F.1 Convective derivative of $\mu_{\zeta|z}$
Consider the Beltrami flow $\zeta^{\prime}=\zeta+{\dot{\zeta}}dt$ under dynamics of the adjacency metric; $\bar{\partial}_{\zeta}{\dot{\zeta}}=\partial_{t}\mu_{a}$. We would like to calculate how $\mu_{\zeta|z}$ changes under this flow while holding the tension metric $\mathbf{g}$ fixed. To find the new Beltrami coefficient $\mu_{\zeta^{\prime}|z}$, we need to compose the maps $\zeta^{\prime}\mapsto\zeta$ and $\zeta\mapsto z$. Drawing the map compositions as a diagram
[…]
Beltrami flow induces $\mu_{\zeta|\zeta^{\prime}}=\bar{\partial}_{\zeta}{\dot{\zeta}}dt/(1+\partial_{\zeta}{\dot{\zeta}}dt)\approx\bar{\partial}_{\zeta}{\dot{\zeta}}dt=\partial_{t}\mu_{a}dt$. To find the Beltrami coefficient of the inverse of a map, one applies the chain rule to obtain the relation
$\mu_{\zeta^{\prime}|\zeta}=-\frac{\overline{\partial_{\zeta}\zeta^{\prime}}}{\partial_{\zeta}\zeta^{\prime}}\mu_{\zeta|\zeta^{\prime}}.$ (74)
Now composing with $\zeta\mapsto z$ gives
$\mu_{\zeta^{\prime}|z}$ $(\zeta^{\prime},\bar{\zeta}^{\prime})=\frac{\mu_{\zeta^{\prime}|\zeta}+\mu_{\zeta|z}\,e^{-i2\phi_{\zeta^{\prime}|\zeta}}}{1+\bar{\mu}_{\zeta^{\prime}|\zeta}\,\mu_{\zeta|z}\,e^{-i2\phi_{\zeta^{\prime}|\zeta}}}$ (75)
$\approx\frac{-\bar{\partial}_{\zeta}{\dot{\zeta}}dt+\mu_{\zeta|z}(1+i2\mathrm{Im}[\partial_{\zeta}{\dot{\zeta}}]dt)}{1-\mu_{\zeta|z}\partial_{\zeta}\bar{v}_{(\zeta)}dt}$
$\approx\mu_{\zeta|z}-\left(\partial_{t}\mu_{a}-i\mu_{\zeta|z}\mathrm{Im}[\partial_{\zeta}{\dot{\zeta}}]+\mu_{\zeta|z}^{2}\partial_{t}\bar{\mu}_{a}\right)dt,$
where the phase factor is $e^{-i2\phi_{\zeta^{\prime}|\zeta}}=(1-\overline{\partial_{\zeta}{\dot{\zeta}}})/(1-\partial_{\zeta}{\dot{\zeta}})\approx 1+i2\mathrm{Im}[\partial_{\zeta}{\dot{\zeta}}]dt$. The displacement change from $\zeta$ to $\zeta^{\prime}$ leads to an additional advective term $\mathrm{Re}[\bar{\nu}\partial_{\zeta}]\mu_{z}dt$. Taken together, one arrives at
$\frac{d}{dt}\mu_{\zeta|z}$ $=\frac{\mu_{\zeta^{\prime}|z}(\zeta^{\prime},\bar{\zeta}^{\prime})-\mu_{\zeta|z}(\zeta,\bar{\zeta})}{dt}$ (76)
$=\partial_{t}\mu_{\zeta|z}-\partial_{t}\mu_{a}+\mathrm{Re}[\bar{v}_{(\zeta)}\partial_{\zeta}]\mu_{\zeta|z}$
$\quad+i2\mu_{z}\mathrm{Im}[\partial_{\zeta}{\dot{\zeta}}]-\mu_{\zeta|z}^{2}\partial_{t}\bar{\mu}_{a}$ (77)
The final, nonlinear term can be neglected for weak anisotropy $|\mu_{\zeta|z}|\ll 1$ and captures strain-induced rotation. Figure 5 provides a geometric explanation of Eq. (76). The term $\partial_{t}\mu_{\zeta|z}$ arises from the intrinsic dynamics of $z$, in addition to the coordinate transformation $\zeta\mapsto\zeta^{\prime}$.
Figure 5: Illustration of the geometry of Eq. (76). The adjacency metric represents a triangulation where each edge has length 1, and therefore undergoes a shear deformation if adjacency is changed by an edge flip (T1 transition). The tension metric remains fixed under a T1. As a result, the tension triangle shape acquires anisotropy.
F.2 Variation of $|\mu_{\zeta|w}|^{2}$
In the main text, we argue that T1s relax the “total” Beltrami coefficient $\mu_{\zeta|w}$ of the map from $\bf a$}-isothermal coordinates $\zeta$ to the physical state $w$. We therefore need to calculate the variation of $\mu_{\zeta|w}$ under a small variation of the adjacency metric ${\bf a}(\bm{\zeta})\mapsto{\bf a}^{\prime}(\bm{\zeta})={\bf a}(\bm{\zeta})+\delta\mathbf{a}(\bm{\zeta})$. First, note that $\mu_{\zeta|w}$ is independent of $\mathbf{g}$ since the map $\zeta\mapsto w$ bypasses the intermediate (reference) configuration defined by $\mathbf{g}$ [cf. (46)]. Now let $\zeta^{\prime}$ be the $a^{\prime}$-isothermal coordinates, which must satisfy the Beltrami equation $\bar{\partial}_{\zeta}\zeta^{\prime}=\mu_{\delta a}$. Drawing the map compositions as a diagram
[…]
makes clear that the Beltrami coefficient of the new map $\zeta^{\prime}\mapsto w$ is calculated by composing $\zeta^{\prime}\mapsto\zeta\mapsto w$ [red path in Eq. (78)]. We first use that $\zeta\mapsto\zeta^{\prime}$ is simply the Beltrami flow induced by $\delta\mathbf{a}$ so $\mu_{\zeta|\zeta^{\prime}}=\bar{\partial}_{\zeta}\zeta^{\prime}/\partial_{\zeta}\zeta^{\prime}=\delta\mu_{a}$. The inverse is obtained using Eq. (74), giving
$\mu_{\zeta^{\prime}|\zeta}=-\frac{\overline{\partial_{\zeta^{\prime}}\zeta}}{\partial_{\zeta^{\prime}}\zeta}\mu_{\zeta|\zeta^{\prime}}=-\frac{\overline{\partial_{\zeta^{\prime}}\zeta}}{\partial_{\zeta^{\prime}}\zeta}\delta\mu_{a}$ (79)
Since $|\mu_{\zeta^{\prime}|\zeta}|\ll 1$, the Beltrami coefficient of $\zeta^{\prime}\mapsto\zeta\mapsto w$ can be calculated by linearizing the Beltrami composition formula Eq. (48):
$\mu_{\zeta^{\prime}|w}$ $\approx\mu_{\zeta^{\prime}|\zeta}+\frac{\overline{\partial_{\zeta^{\prime}}\zeta}}{\partial_{\zeta^{\prime}}\zeta}\mu_{\zeta|w}=\frac{\overline{\partial_{\zeta^{\prime}}\zeta}}{\partial_{\zeta^{\prime}}\zeta}\left(\mu_{\zeta|w}-\delta\mu_{a}\right)$ (80)
And therefore, we conclude:
$|\mu_{\zeta^{\prime}|w}|^{2}$ $=|\mu_{\zeta|w}|^{2}+(\mu_{\zeta|w}\delta\bar{\mu}_{a}+\bar{\mu}_{\zeta|w}\delta\mu_{a})$ (81)
$\Rightarrow\frac{\delta|\mu_{\zeta|w}|^{2}}{\delta\bar{\mu}_{a}}$ $=-\mu_{\zeta|w}.$ (82)
Appendix G Examples
Here, we provide three concrete examples to illustrate the continuum equations derived in the main text. First, we consider two static examples. Since the static problem is mathematically equivalent to linear elasticity, these examples have no novel content, but serve to clarify the logic of the mapping from the tension manifold to the physical state. Second, we illustrate the dynamics equations by considering the example of positive tension feedback, building on the mean-field analysis in Ref. [30].
G.1 Statics: flat tension metric
As a first example, we take a uniform, incompressible, rectangular tissue patch. The set-up is sketched in Fig. 6. Let the adjacency metric be the Euclidean metric, corresponding to uniformly distributed isotropic cells. The $a$-isothermal $\zeta$ coordinates are simply Cartesian coordinates in the unit square $0\leq\zeta_{1},\zeta_{2}\leq 1$. In this square shape, the cell density is uniform, and cell shapes are isotropic. Next, we need to specify the active tension configuration, i.e., the tension metric $g$. In this example, we consider an anisotropic metric, $g_{\zeta\zeta}=2,\quad g_{\zeta\bar{\zeta}}=s$ where the real parameter $0<s\ll 1$ determines the strength of the anisotropy. In Cartesian notation, $\mathbf{g}(\zeta_{1},\zeta_{2})=\mathrm{diag}[1+s,1-s]$. Microscopically, this tension metric means that vertical cell-cell interfaces are under higher tension than horizontal ones, since the tension metric encodes the tension on interfaces transverse to a line element $d\mathbf{r}$ (cf. Fig. 2). The Beltrami coefficient reads $\mu_{g}=s/(2+\sqrt{(1+s)(1-s)})\approx s/4$, where the approximation is for $s\ll 1$. To remove this anisotropy, the $\mathbf{g}$-isothermal coordinates need to create a corresponding shear, $z(\zeta,\bar{\zeta})=\zeta+\mu_{g}\bar{\zeta}$. This fulfills the Beltrami Eq. (10). In Cartesian notation, $\mathbf{z}=\mathrm{diag}[1+\mu_{g},1-\mu_{g}]\cdot\bm{\zeta}$.
Figure 6: Elementary example for tension- and adjacency metrics.
Since $\det\mathbf{g}\approx 1$, tension and cell densities are (approximately) equal in this example. Further, since $\mathbf{g}$ is constant, it has no curvature, $K_{g}=0$: the $z$-configuration has constant cell density and constant pressure $p_{0}$. It is therefore mechanically balanced in the absence of external forces. In other words, $u_{\alpha}=0$ solves the linear elasticity force-balance equation Eq. (33). This elementary example illustrates how tension anisotropy controls tissue shape.
Let us now consider applying a constant boundary traction force $b_{1}=0,b_{2}=\pm b$ to the upper (++) and lower (-) edge of the rectangular tissue patch, (the left and right sides remain free). This results in a potential displacement $u_{\alpha}=f_{\alpha}+\partial_{\alpha}\theta$ from the tension reference $z$ to the physical configuration $w=z+u(z,\bar{z})$. By Eq. (36), the potential $\theta$ must be biharmonic, $\Delta^{2}\theta=0$. The traction forces $\sigma\cdot\mathbf{n}_{\pm}=[0,\pm b]$ at the top and bottom edges require $\sigma_{22}|_{\pm}=p_{0}(\partial_{1}^{2}\theta)|_{\pm}=b$. We can hence take $\theta=b\operatorname{Re}[z]^{2}/(2p_{0})$, which implies a uniform stress $\bm{\sigma}=\mathrm{diag}\,(0,b)$ in the bulk and, thus, also fulfills the free BCs on the left and right edge. This result is expected by simple physical reasoning. To ensure that the displacement $u_{\alpha}$ is incompressible, we add a conformal contribution f(z)=bzf(z)=-bz (in this case, simply a uniform scaling).
G.2 Statics: non-flat tension metric
Next, we consider an example where the tension network has curvature. We consider a circular tissue patch. The adjacency metric is the Euclidean metric in the unit disk. The $\mathbf{a}$-isothermal $\zeta$-coordinates are simply Cartesian coordinates. In $\zeta$-coordinates, the cell density is uniform. We now define the tension metric in the disk configuration:
$(g_{\alpha\beta})=\begin{pmatrix}1+\zeta_{1}^{2}/h^{2}&\zeta_{1}\zeta_{2}/h^{2}\\ \zeta_{1}\zeta_{2}/h^{2}&1+\zeta_{2}^{2}/h^{2}\end{pmatrix},\;h(\zeta,\bar{\zeta})^{2}=1-\zeta\bar{\zeta}$ (83)
We use round brackets to highlight that this matrix uses Cartesian, not complexified, coordinates. It is the metric of a unit hemisphere, a height $h$ above the unit circle in Monge parametrization (Fig. 7). Since $\det\mathbf{g}=1/h(\zeta,\bar{\zeta})^{2}$, the “tension density” is higher at the margin than at the center.
We now need to find the transformation $z(\zeta,\bar{\zeta})$ that makes this tension metric isotropic. Because of rotation symmetry, we expect a radial dilation. Indeed, the $g$-isothermal coordinates are given by the stereographic projection [40] of the hemisphere, which is a conformal map:
$z(\zeta,\bar{\zeta})=\frac{\zeta}{1+h(\zeta,\bar{\zeta})}$ (84)
The conformal factor reads $\lambda_{g}(z,\bar{z})=2/(1+z\bar{z})$, which fulfills $\lambda_{g}=1$ on the boundary. By Eq. (12), the curvature is constant, $K_{g}=1$, as expected for a unit sphere.
Due to curvature, area distortion in the isothermal coordinates, and thus cell density gradients in the tension reference state $z$ are unavoidable. For an incompressible tissue, these must be compensated by an isogonal displacement $w-z=u=f(z)+2\bar{\partial}_{z}\theta$ (Eq. (33)). Incompressibility means $(\det\mathbf{D}w)^{-1}\sqrt{\det g(z,\bar{z})}=1$ where $(Dw)_{\alpha\beta}=\partial_{\alpha}w_{\beta}$ is the Jacobian. Expanding $\det\mathbf{D}w=\det[\delta_{\alpha\beta}+\partial_{\alpha}u_{\beta}]$ with $w_{\alpha}=f_{\alpha}+\partial_{\alpha}\theta$ to linear order, this implies
$\Delta\theta+\partial_{\alpha}f_{\alpha}=\frac{2}{z\bar{z}-1}$ (85)
We now obtain a rotation-symmetric solution to Eq. (85) as a function of $r=\sqrt{z\bar{z}}$. First, the only rotation-symmetric conformal mode is a uniform scaling $f(z)=Cz$. Further, we fix an irrelevant global constant by setting $\theta(0)=0$. The equation for $\theta(r)$ then reads
$\Delta\theta(r)+C=\frac{2}{1+r^{2}}$ (86)
with BCs $\theta(0)=0$, $\partial_{r}\theta(r=0)=0$, and $\partial_{r}\theta(r=0)=0$. The last condition follows from $\sigma\cdot\mathbf{n}=0$ (vanishing traction forces at the boundary). The strain $C=\partial_{\alpha}f_{\alpha}$ of the conformal map is chosen to fulfill the last BC. Eq. (86) is solved by $\theta(r)=\tfrac{1}{4}[-r^{2}+2\log(1+r^{2})]$ with $C=1$.
Within the disk, $\sigma_{\alpha\beta}^{\mathrm{tot}}=-p_{0}\epsilon_{ac}\epsilon_{bd}\partial_{\gamma}\partial_{\delta}\theta\neq 0$ (Eq. (35)). These internal residual stresses, shown in Fig. 7, result from curvature of the tension manifold. Importantly, such stresses will also be present for finite cell compressibility. In the limit of fully compressible cells, no force-balanced configuration exists because sustaining the pressure field implied by the tension manifold curvature requires a finite cell compressibility.
Figure 7: A half-sphere tension manifold induces a non-constant pressure field and a non-vanishing total stress even for traction-free boundary conditions. Note that the total stress is isotropic in the center and purely tangential along the boundary, as is required by the traction-free boundary conditions.
G.3 Dynamics: positive tension feedback
We now consider an example of tissue flow driven by tension dynamics. We translate the results of Ref. [30], which analyzed a model for positive tension feedback with mean-field techniques, into the notation of the present manuscript. For simplicity, we consider the case where $w\approx z$, i.e., no tension-cell density incompatibility or external forces.
In the mean-field setting, a tissue patch is described by the local distribution of tension triangles (correlations between adjacent triangles are neglected), each of which is characterized by a Beltrami coefficient $\mu_{\zeta|z}^{\triangle}$. The macroscopic order parameters are moments of this distribution. The mean $\langle\mu_{\zeta|z}^{\triangle}\rangle=\mu_{\zeta|z}$ defines the macroscopic tension anisotropy, while the second moment $\langle|\mu_{\zeta|z}^{\triangle}|^{2}\rangle=q^{2}$ measures how anisotropic individual triangles are. Note that $q>|\mu_{\zeta|z}|$: e.g., in a disordered triangulation, triangles can be anisotropic with no net tissue-scale anisotropy. (In the notation of Ref. [30], $\mathbf{N}=\mu_{\zeta|z}/q$, neglecting correlations between $\arg\mu_{\zeta|z}^{\triangle}$ and $|\mu_{\zeta|z}^{\triangle}|$). In general, a wider distribution of tension anisotropies leads to a higher T1 rate, since there are more triangles “close” to the T1 threshold.
The parameter $q$ determines the T1 rate $\gamma_{\mathrm{T1}}(q)$: the higher $q$, the higher the rate of T1s, with $\gamma_{\mathrm{T1}}\mapsto\infty$ as $q\mapsto 1$. The dynamics of the tension anisotropy $\mu_{\zeta|z}$ combined two effects: amplification due to positive feedback with a characteristic timescale $t_{\tau}$ and degradation of anisotropy by T1s:
$\partial_{t}\mu_{\zeta|z}=\frac{1}{t_{\tau}}\mu_{\zeta|z}-\alpha(q)\gamma_{\mathrm{T1}}(q)\mu_{\zeta|z}.$ (87)
The coefficient $\alpha$ can be calculated from the mean field theory. One finds that the combined coefficient $1-\alpha\gamma_{\mathrm{T1}}t_{\tau}<0$ at long times, so that the dynamics eventually erase any tension anisotropy. However, the transient remodeling of the adjacency metric, due to the Beltrami flow $\bar{\partial}_{\xi}(\partial_{t}\zeta)=\gamma_{\mathrm{T1}}\mu_{z}$, leads to a net plastic tissue deformation due to the initial tension anisotropy.